### v1

# New black hole mergers in the LIGO–Virgo O3 data from a gravitational wave search including higher-order harmonics

Abstract
Nearly all of the previous gravitational wave (GW) searches in the LIGO–Virgo data included GW waveforms with only the dominant quadrupole harmonic, i.e., omitting higher-order harmonics which are predicted by general relativity. Based on the techniques developed in Wadekar et al. (2023a, b), we improve the IAS pipeline by () introducing higher harmonics in the GW templates, () downweighting noise transients (“glitches”) to improve the search sensitivity to high-mass and high-redshift binary black hole (BBH) mergers. We find 14 new candidate BBH mergers with on running our pipeline over the public LIGO–Virgo data from the O3 run (we use the detection threshold as the astrophysical probability, , being over 0.5, following the approach of other pipelines). We also broadly recover the high-significance events from earlier catalogs, except some which were either vetoed or fell below our SNR threshold for trigger collection.
[…]
I Introduction
Gravitational waves (GW) from roughly 100 binary black hole (BBH) mergers have been detected in the data from the first three observing runs of the LIGO–Virgo–Kagra collaboration (LVK), combining detections by the LVK Abbott et al. (2016a, 2019, 2021); The LIGO Scientific Collaboration et al. (2021); Abbott et al. (2021) and by other independent pipelines Venumadhav et al. (2019, 2020); Olsen et al. (2022); Nitz et al. (2019, 2020, 2021, 2021); Chia et al. (2023) who access the data made publicly available through the Gravitational Wave Open Science Center (GWOSC) Vallisneri et al. (2015). Nearly all the previous template-bank based searches have approximated GW waveforms to have only the dominant quadrupole mode, , where the frequency of GWs is twice the orbital frequency.
The mode appears at the lowest order in post-Newtonian (PN) expansion, but there are a number of higher harmonics originating from the subsequent orders of the PN expansion (e.g., the mode has thrice the orbital frequency) Thorne (1980); Pan et al. (2011). Approximate formulae for the relative strength of frequency-domain amplitudes of two of the higher-order modes (HM) are given by Wadekar et al. (2023a)
where is the mass ratio (), is the orbital frequency and is the binary inclination ( corresponds to the edge-on configuration). These harmonics have a number of important applications for characterizing properties of BBHs. If one uses the mode waveform alone in inferring the BBH properties, we have the following important degeneracies between different binary parameters: luminosity distance and ; and effective spin (); polarization and initial orbital phase (). All three degeneracies are broken by HM Mills and Fairhurst (2021); Varma et al. (2014). This follows from the amplitudes of HM having a different dependence on and as compared to that of the mode (see Eq. (1)), and similarly the phases of HM have different dependence on . It is worth mentioning that the degeneracy between distance and inclination is currently one of the most important limiting factors in inferring the Hubble parameter from GW systems and HM can be crucial in breaking this degeneracy. Similarly, events which have HM also have the best measured mass ratios (e.g., Abbott et al. (2020a, b)). HM become important in the strong gravity regime close to merger and are thus useful to test the multipolar structure of gravity Kastha et al. (2018, 2019); Dhanpal et al. (2019); Islam et al. (2020); Capano et al. (2021), and to study properties of the merger remnant such as the recoil kick.
Multiple waveform models including HM have been recently made available London et al. (2018); García-Quirós et al. (2020); Mehta et al. (2017), and they are currently being used ubiquitously in parameter estimation studies. However, HM are still not included in current search pipelines (as of now, the only exception is Ref. Chandra et al. (2022), which searched for binaries with nearly edge-on inclinations). The reason behind this is that for making template banks with the mode, one only needs to sample over the masses and spins of black holes (). However, for constructing banks including HM, one needs to additionally sample over the inclination and initial phase (, ). These additional degrees of freedom can cause: the size of the template banks and the cost of search to increase significantly (by a factor of 100 for the case of HM Harry et al. (2018); Chandra et al. (2022); Schmidt et al. (2023)), loss of sensitivity in the search as there is increased number of background triggers Capano et al. (2014). To tackle the first issue, we found more efficient ways of including HM in our template banks (which we presented in detail in our first companion paper Wadekar et al. (2023a)). Regarding the second issue, we devise new detection statistics which includes marginalization over the amplitudes and phases of higher modes in our second companion paper Wadekar et al. (2023b). We present an overview of these new techniques in Section II.
We see from Eq. (1) that, at a given orbital frequency, the relative HM amplitudes increase as a power-law with . Moreover, due to the shape of the current detector power spectral density (PSD), the signal-to-noise ratio (SNR) contribution of HM is additionally boosted at high masses and redshifts. In this regime, the mode can cut off below the detector’s sensitive band (the detector PSD sharply increases below Hz) and HM, being at higher frequencies, naturally have an increased contribution to the SNR. The background corresponding to the high-mass regime is however heavily dominated by short-duration noise transients (called instrumental “glitches”). This is because the high-mass waveforms have a very short duration in band, which makes it easy for transients like blip glitches to mimic real signals. We therefore introduce additional methods in the IAS pipeline to downweight the background at high masses. These methods are detailed in our second companion paper Wadekar et al. (2023b) and we present an overview in Section II.3.
Running our new search pipeline over the public LIGO–Virgo data from the third observing run (O3), we found new candidate events with properties shown in Figs. 1 and 2. We further discuss these new events and also show our search results for the events from the GWTC-3 and 4-OGC catalogs in Section III. We review the interesting astrophysical properties of the new candidate events in Section IV. We analyze the impact of our astrophysical prior and comment upon the gain/loss of sensitivity due to adding HM in Section V, and we conclude in Section VI.
II Pipeline improvements
Here, we present an outline of the additions that we made to the -only IAS pipeline Olsen et al. (2022); Venumadhav et al. (2019).
II.1 Adding HM in template banks
In this sub-section, we summarize the methodology for constructing template banks for HM from our companion paper: Wadekar et al. (2023a). For each mode template, we use a combination of post-Newtonian formulae and machine learning tools to model normalized aligned-spin templates corresponding to the and modes. We then filter the data separately with the three templates and store the complex SNR timeseries , see figure 1 of Wadekar et al. (2023a). As a result, our HM search matched-filtering cost is just that of a -only search (in contrast to being as in previous HM search methods, where the different modes are combined and templates sample over inclination and initial orbital phase Chandra et al. (2022); Schmidt et al. (2023); Harry et al. (2018)).
We divide our search space into 17 template banks, which are divided according to the shape of normalized waveforms corresponding to the binary parameters (the shape of normalized waveforms is roughly related to the binary merger frequency). The parameter space corresponding to each bank is shown in the bottom panel of Fig. 7 in the Appendix. Some of the banks are further divided into subbanks based on the range of chirp masses they cover. Note that the division into banks and subbanks is only to facilitate covering the parameter space with templates of the form that we use, and the final ranking statistic we assign to each trigger is weighed against those of all other triggers (over the entire parameter space) to estimate its global significance. Other than adding HM, we also improved our template banks by using a machine learning tool called random forest regressor (which helps us reduce the dimensionality of our template banks and enables us to search with 3040% fewer templates than before) Wadekar et al. (2023a).
II.2 New statistics for combining SNR timeseries from different harmonics
We present the techniques for combining the timeseries in our companion paper (Wadekar et al. (2023b)) and here we provide a brief overview. Combining SNR timeseries of different modes incorrectly can lead to a large loss of sensitivity of the pipeline Capano et al. (2014). We use the Neyman–Pearson lemma to develop an optimal statistic to combine the SNR timeseries for different harmonics; our detection statistic is the ratio of the Bayesian evidence under the signal and the noise hypothesis. Ideally, one would marginalize over all the intrinsic and extrinsic binary parameters to calculate the evidence under the signal hypothesis. We develop two different approximate versions of the statistic to be used in different parts of our pipeline as detailed in the sub-sections below.
II.2.1 Single-detector marginalized statistic
We first developed an approximate statistic to collect individual detector triggers. Similar to other pipelines, we set a particular SNR threshold () to collect triggers from individual detectors. The primary reason behind this is to reduce the number of triggers passed to the next steps of the pipeline (coincidence analysis, veto and coherent score calculation) as these steps are computationally expensive (another reason is storage space constraints when saving a large number of triggers on disk). For collecting the triggers in a -only search, one can simply use . For the HM case, if we naively use , the number of collected triggers increases by a factor of over (using the same value of as the -only case Wadekar et al. (2023b)). To ameliorate this, we develop in Ref. Wadekar et al. (2023b) a new marginalized statistic () which exploits the fact that the values of and are bounded for physical signals (as opposed to Gaussian noise triggers where the values and can be arbitrarily large). The statistic includes marginalization over physically possible amplitude ratios of higher harmonics. To make the computation of faster, we maximize (instead of marginalizing) over the phases of HM and also over all the remaining intrinsic and extrinsic binary parameters (although this can be improved in future studies). We collect single-detector triggers using and we use a larger value of the collection threshold as compared to the -only case. The downside of increasing the value of is that low-SNR events can fall below our threshold bar. This is indeed the reason behind our pipeline missing some of the low-SNR events reported by -only pipelines in the previous literature, see Section III.3 below for more details. The particular choice of that we made in the current paper is currently not a bottleneck for our search and can be improved in the future.
II.2.2 Coherent marginalized statistic
Once we collect single detector triggers, we check if the same template(s) fired in other detectors at a similar time. To rank the resulting coincident triggers (for calculating their false-alarm rate (FAR) and subsequently ), we develop a more accurate statistic which also includes coherence between different detectors for ranking the events. For example, we exploit the fact that astrophysical signals have correlated phase and time differences between the detectors, while the noise triggers have random distributions Nitz et al. (2017). In the previous sub-section, we maximized over extrinsic binary parameters, but marginalizing over them is the more optimal strategy. Our coherent ranking statistic thus includes marginalization over extrinsic parameters in a similar way to the IAS -only search methodology given in Appendix D of Olsen et al. (2022) (hereafter O22) except generalized to account for higher modes. In addition, we also include marginalization over the amplitude of HM (which effectively corresponds to marginalizing over mass ratios); the only parameters that we currently do not marginalize over are the intrinsic binary parameters corresponding to the 22 waveform, as we use only the best-fit 22 template in our calculations. Details regarding our coherent statistic will be given in upcoming papers Wadekar et al. (2023b); Roulet et al. (2023).
II.3 Removing/downweighting noise transients
In addition to the changes involving higher modes, we also improve the sensitivity of our search for high masses compared to previous IAS pipelines by developing new tools. We also discuss in this sub-section methods to tailor the background reduction techniques in our -only pipeline for the case of HM.
II.3.1 Band eraser tool for data preprocessing
In our -only pipeline, in a data pre-processing step (before we matched-filter with templates), we identify regions of bad data segments which have excess power localized to particular frequency ranges and timescales (see section C of Venumadhav et al. (2019)). We then remove the data in the flagged regions (i.e., make “holes”) and then fill the holes using an “inpainting filter” (see figure 6 of Zackay et al. (2021)). Note that the holes remove the entire data from a particular time chunk. Sometimes, however, the noise transients are only localized to particular frequency bands (particularly the low-frequency end 100 Hz) and do not affect astrophysical signals which dominantly have power in a different frequency band.
Therefore, before making holes in entire time-chunks, we try to smoothly remove bad time-frequency regions (or bands) in the spectrogram. For this, we develop a new tool we call band eraser. We divide the spectrogram into multiple bands (each with dimensions ). We further split each band into multiple chunks (of ) and measure the number of chunks which have power above a certain threshold. If this number is above a certain threshold, we remove the entire band from our analysis. We calculate this threshold using Poisson statistics, making sure that, on average, in Gaussian noise, one band is removed per s (we also add an additional safety factor to the threshold to avoid removing gravitational wave signals).
We also noticed that some of the noise transients had power spread over long timescales ( s) and thus were sometimes missed by the localized excess-power tests (with timescales s) that we had used earlier to identify regions for making holes (see Table II of Venumadhav et al. (2019)). Empirically, we noticed that such disturbances predominantly pollute the low-frequency ( Hz) end, and therefore particularly reduced the sensitivity of our pipeline to high-mass mergers. We find that using the band eraser also helped in mitigating some such disturbances. We leave further discussion of the band eraser tool to Ref. Wadekar et al. (2023b).
II.3.2 Veto procedure
The -only IAS pipeline included a series of signal consistency checks in order to veto triggers that are likely to arise from noise transients, in order to improve the search sensitivity (see section I of Venumadhav et al. (2019)). On the coincident triggers, we perform similar veto procedures as the -only pipeline, except for the following change. For a particular trigger, we first subtract our best-fit and waveforms from data and then run the same signal consistency tests as earlier on the residual data. In the limit in which the higher harmonics are nearly orthogonal to the waveform, the consistency tests can be left largely unchanged. We aim to modify these tests to also check the consistency of the and parts of the signals in the future.
In previous quadrupole-only searches, we reduced the noise background by accounting for the non-stationarity of the PSD Zackay et al. (2021). In the procedure, we match-filtered a reference template from the bank with the data and registered its SNR . We then estimated the variation in PSD as a function of time by averaging the power within rolling windows of length s. In this search, we calculate the correction for the PSD drift in the same way as before by only using waveforms from our template bank (and not including HM waveforms). In the limit in which the fluctuations of the PSD are described by a scalar over all frequencies, this procedure returns an unbiased estimate of the drift correction.
II.3.3 Ranking statistic
Even after performing the above steps, there are still some noise triggers which survive. We use the ranking statistic as a final tool to downweight such triggers. Beyond upweighting triggers which have good coherence among multiple detectors, an important job of the ranking statistic is to also downweight triggers based on the probability with which similar triggers occur in the background. We simulate 2000 instances of background runs using the method of timeslides (where we shift one of the detectors by unphysical amounts which are more than the light crossing time between the detectors).
In the -only case, for a trigger with SNR associated with a template , this probability under the noise hypothesis is . In the case of Gaussian noise, follows a (i.e., ). However, we empirically measure by constructing a histogram of the values of background triggers associated with the template . In our previous papers, we labelled this term as the “rank function” (), see section J of Ref. Venumadhav et al. (2019).
In a similar way for the HM case, if the noise were Gaussian, separately follow distributions (note that we orthogonalize templates of different harmonics), and their sum thus follows a distribution with 6 d.o.f. This motivated us to make the approximation and again measure this probability distribution empirically by constructing a 1D histogram of values of background triggers associated with the template . We plan to explore beyond the current assumption in a future study.
Note that in the previous IAS studies, is assumed to be the same for all templates in a subbank. However, especially for high masses, this term can have different values even for different templates within a subbank, see e.g., Abbott et al. (2016b); Nitz et al. (2017); Wadekar et al. (2023b). In order to optimize our analysis, we follow the technique in Ref. Wadekar et al. (2023b), where we first cluster the templates within each sub-bank based on their sensitivity to glitches (as quantified empirically by the relative fraction of loud and faint triggers belonging to each template). We then separately construct the histograms of background triggers for each template group to estimate the noise probabilities of triggers associated to templates within that group.
In a future paper, we plan to quantify the improvement in volume-time sensitivity due to the background reduction procedures outlined in this section, by injecting simulated gravitational-wave signals in the data and comparing the fraction of injections recovered by our pipeline.
Sr. | New candidate | Bank | PE (90% credibility) | (evidence) | IFAR (yr) | |||||||||
No. | event | HM | HM+Prec. | per bank | overall | |||||||||
1 | GW190605_025957 | 13,0 | ||||||||||||
2 | GW190806_033721 | 8,1 | ||||||||||||
3 | GW190524_134109 | 10,0 | ||||||||||||
4 | GW191113_103541 | 9,0 | ||||||||||||
5 | GW190615_030234 | 10,0 | ||||||||||||
6 | GW190604_103812 | 12,0 | ||||||||||||
7 | GW191228_195619 | 14,0 | ||||||||||||
8 | GW200304_172806 | 8,1 | ||||||||||||
9 | GW190530_030659 | 4,2 | ||||||||||||
10 | GW190511_163209 | 8,0 | ||||||||||||
11 | GW200210_100022 | 9,0 | ||||||||||||
12 | GW200301_211019 | 1,2 | ||||||||||||
13 | GW190911_195101 | 12,0 | ||||||||||||
14 | GW190530_133833 | 12,0 |
Event | Bank | PE (90% credibility) | (evidence) | IFAR (yr) | |||||||||
HM | HM+Prec. | per bank | overall | ||||||||||
GW190711_030756 | 7,1 | ||||||||||||
GW190707_083226 | 7,1 | ||||||||||||
GW200109_195634 | 10,0 | ||||||||||||
GW190818_232544 | 8,1 | ||||||||||||
GW190906_054335 | 7,1 |
GWTC-3 events | Bank | IFAR (yr) | ||||||||
/comment | IAS | GWTC-3 | 4-OGC | IAS | GWTC-3 | 4-OGC | ||||
GW190403_051519 | BBH_9 | |||||||||
GW190408_181802 | BBH_2 | |||||||||
GW190412_053044 | BBH_3 | |||||||||
GW190413_052954 | BBH_6 | |||||||||
GW190413_134308 | BBH_7 | |||||||||
GW190421_213856 | BBH_7 | |||||||||
GW190426_190642 | Veto (BBH_12) | 25.82 | 45.46 | |||||||
GW190503_185404 | BBH_4 | |||||||||
GW190512_180714 | BBH_1 | |||||||||
GW190513_205428 | BBH_4 | |||||||||
GW190514_065416 | BBH_7 | |||||||||
GW190517_055101 | BBH_2 | |||||||||
GW190519_153544 | BBH_7 | |||||||||
GW190521_030229 | Veto (BBH_16) | 109.23 | 180.37 | |||||||
GW190521_074359 | BBH_6 | |||||||||
GW190527_092055 | BBH_6 | |||||||||
GW190602_175927 | BBH_9 | |||||||||
GW190701_203306 | Veto (BBH_7) | 34.87 | 55.9 | |||||||
GW190706_222641 | BBH_8 | |||||||||
GW190707_093326 | BBH_0 | |||||||||
GW190719_215514 | BBH_8 | |||||||||
GW190720_000836 | BBH_0 | |||||||||
GW190725_174728 | Veto (BBH_3) | 27.35 | 55.99 | |||||||
GW190727_060333 | BBH_6 | |||||||||
GW190728_064510 | BBH_0 | |||||||||
GW190731_140936 | BBH_6 | |||||||||
GW190803_022701 | BBH_6 | |||||||||
GW190805_211137 | BBH_6 | |||||||||
GW190828_063405 | BBH_4 | |||||||||
GW190828_065509 | Veto (BBH_1) | 50.14 | 56.48 | |||||||
GW190909_114149111GW190909_114149 is a marginal candidate in GWTC-3 which has from our pipeline. | BBH_7 | |||||||||
GW190915_235702 | BBH_4 | |||||||||
GW190916_200658 | BBH_8 | |||||||||
GW190917_114630 | Below | |||||||||
GW190924_021846 | BBH_0 | |||||||||
GW190926_050336 | BBH_6 | |||||||||
GW190929_012149 | BBH_9 | |||||||||
GW190930_133541 | Veto (BBH_1) | 36.02 | 45.98 | |||||||
GW191103_012549 | BBH_0 | |||||||||
GW191105_143521 | BBH_0 | |||||||||
GW191109_010717 | Veto (BBH_8) | 82.74 | 164.37 |
GWTC-3 events | Bank | IFAR (yr) | ||||||||
/comment | IAS | GWTC-3 | 4-OGC | IAS | GWTC-3 | 4-OGC | ||||
GW191113_071753 | Below | |||||||||
GW191126_115259 | BBH_0 | |||||||||
GW191127_050227 | BBH_8 | |||||||||
GW191129_134029 | BBH_1 | |||||||||
GW191204_110529 | BBH_2 | |||||||||
GW191204_171526 | BBH_0 | |||||||||
GW191215_223052 | BBH_2 | |||||||||
GW191222_033537 | BBH_7 | |||||||||
GW191230_180458 | BBH_8 | |||||||||
GW200128_022011 | BBH_6 | |||||||||
GW200129_065458 | Veto (BBH_4) | 203.73 | 342.75 | |||||||
GW200202_154313 | BBH_1 | |||||||||
GW200208_130117 | BBH_6 | |||||||||
GW200208_222617 | BBH_8 | |||||||||
GW200209_085452 | BBH_4 | |||||||||
GW200210_092254 | Veto (BBH_3) | 25.53 | 48.5 | |||||||
GW200216_220804 | BBH_8 | |||||||||
GW200219_094415 | BBH_6 | |||||||||
GW200220_061928 | Below | |||||||||
GW200220_124850 | BBH_6 | |||||||||
GW200224_222234 | BBH_6 | |||||||||
GW200225_060421 | BBH_1 | |||||||||
GW200306_093714 | BBH_5 | |||||||||
GW200308_173609 | Below | |||||||||
GW200311_115853 | BBH_4 | |||||||||
GW200316_215756 | BBH_1 | |||||||||
GW200322_091133 | BBH_6 | |||||||||
4-OGC events | Bank | IFAR (yr) | ||||||||
/comment | IAS | GWTC-3 | 4-OGC | IAS | GWTC-3 | 4-OGC | ||||
GW191224_043228 | BBH_1 | 73.0 | ||||||||
GW200106_134123 | BBH_6 | 55.0 | ||||||||
GW200129_114245 | BBH_10 | 63.0 | ||||||||
GW200210_005122 | Below | 70.0 | ||||||||
GW200214_223306 | BBH_8 | 55.0 | ||||||||
GW200305_084739 | Below | 59.0 | ||||||||
GW200318_191337 | BBH_7 | 63.0 |
II.4 Astrophysical prior
In this section, we discuss the astrophysical prior used in our search. This prior is used to assign weights to different templates according to the probability of the template firing under the signal hypothesis : ), and also to downweight triggers with unphysical and ratios as discussed earlier in Section II.2.
We only probe binary black hole (BBH) mergers in our study and leave the search for systems containing neutron stars for future work. The range of our search was restricted to the following space:
where refer to detector-frame (redshifted) masses of the binary. We chose the location of the mass ratio cut for BBHs to restrict to the calibration regime of the IMRPhenomXAS approximant Pratten et al. (2020), while the cut on the individual spins is due to the relativistic Kerr limit.
We picked a simple prior which broadly covers the parameter space (picking a broad prior can help with reweighting the results to a different prior in the future). Our prior on the mass ratio is a uniform one on for ; the effective spin parameter is flat within (with ); the total mass is distributed in a power-law fashion, . Note that we use a prior on the detector-frame masses, and we do not assume a separate prior over redshift .
We drew samples according to this prior (each sample corresponding to the physical parameter set ), and assigned a weight to each sample based on its observable volume. As a proxy for the observable distance, we simply calculated the SNR for the sample at a fixed distance by simulating its waveform using IMRPhenomXHM and using the same reference PSD which was used to make the template banks (weight SNR). Just for reference, for inspiral dominated waveforms, the observable volume roughly scales as Fishbach and Holz (2017). Using these samples, we then utilized the kernel density estimation method to construct probabilities for templates in our template banks (see appendix C of O22 for details on the density estimation method). Fig. 7 in the Appendix shows the cumulative probabilities for our banks, each of which is obtained by adding the weights of all the samples falling within the corresponding bank.
II.4.1 Comparison with astrophysical prior used in our -only banks
It is worth briefly comparing the astrophysical distributions we assumed here to those we assumed in our previous -only O3 searches. We split our banks roughly according to the detector-frame chirp mass (the bank edges were uniform in log ) and further split each bank into subbanks based on the amplitude profiles of normalized waveforms Roulet et al. (2019). We assumed that the total prior astrophysical probability within each of our banks was identical, and that the probability distribution within each bank was flat in the detector-frame component masses and also in the effective spin (without taking the observable distance into account; we take this effect into account in the current search). Note that we also slightly increased our search space in this paper: previously, our banks were limited to and (see figure 2 of O22), while we now use .
III Results
III.1 New candidate events
Using our new search pipeline on the Hanford–Livingston data from the third observing run (O3), we find 14 candidate events with which have not been previously reported in the literature. We list the properties of these new candidates in Table 1. In all tables in this paper, , columns correspond to incoherent SNR obtained from in the particular detector (note that we have orthogonalized the templates of different harmonics, so that they can be added in quadrature). The incoherent values may not correspond to those possible for a physical trigger. Thus, in some cases we see that can even exceed the best-fit parameter estimation (PE) values . Further discussion of the PE runs are given below in Section III.1.1, and we report 90% CL values of some of the binary parameters in Table 1. We show the SNR contribution of individual modes separately in Table 5 in the Appendix.
We also mention the best bank_id, subbank_id for each event; the bank_id roughly increases with mass (see Fig. 7). Note that the high-mass banks in our case have much fewer templates than the low-mass banks (see Table I of Wadekar et al. (2023a)), which leads to considerable heterogeneity in the number of background triggers in the banks. Hence, the values needed for events to be above threshold can be different for different banks.
Other than running the pipeline on coincident triggers, we also artificially simulate background using 2000 timeslides (where we shift one of the detectors by unphysical amounts which are more than the light crossing time between the detectors) and run the pipeline on these in exactly the same way. This allows us to measure the inverse false alarm rate (IFAR) of our coincident triggers. We report this value within each bank and the overall IFAR after combining the background lists from all of our 17 banks. Additionally, for each trigger, the statistic is its probability of being of astrophysical origin; computing it involves estimating the astrophysical rate of events in addition to their IFARs. We estimate the values in a manner that is similar to that outlined in Appendix B of O22, from the distribution of foreground and background triggers according to
where in the above formula corresponds to our final ranking statistic value for the candidate events (see Ref. Wadekar et al. (2023b) for the full expression), and corresponds to number density values from a 1D histogram of the ranking statistic (which we separately calculate for the signal and background cases). Note that as the templates in our banks have astrophysical probabilities associated with them (see Section II.4), and we use these values in calculating . Hence, our IFAR and values are also dependent on our assumption of astrophysical distributions of binary parameters. We expect this dependence to be more significant for the ‘overall’ IFAR in comparison to the ‘per bank’ IFAR values quoted in Table 1. We therefore separately provide the ‘per bank’ values in our table, so one can reweight them in a rough way for different astrophysical distributions than the ones assumed here.
III.1.1 Parameter estimation runs
We perform parameter estimation (PE) runs using the package222https://github.com/jroulet/cogwheel Roulet et al. (2022) with the sampler Lange (2023). We use priors that are uniform in detector-frame constituent masses and comoving volume-time () throughout this paper. Our fiducial prior for the BH spins is flat in effective spin , but we also compare results with those obtained using an isotropic prior on individual spins (which is used for PE in the GWTC-3 and 4-OGC catalogs). Other extrinsic parameters have the standard geometric priors used in GWTC-3 Abbott et al. (2021). We compute likelihoods using the relative binning/heterodyning method described in Ref. Zackay et al. (2018) as implemented in cogwheel. The relation between redshift and luminosity distance is based on the Planck15 cosmology Ade et al. (2016). Our fiducial parameter estimation runs are based on the IMRPhenomXODE waveform approximant Yu et al. (2023) and include the effect of both HM and precession. Table 1 reports marginalized 1D posterior values for a relevant subset of parameters.
Other than the fiducial runs, we also perform two separate PE runs without precession, i.e., restricting to binaries with aligned spins: one run with a -only waveform approximant and the other with an approximant that includes HM as well. This helps us calculate the relative evidence (Bayes factor) due to including the effects of HM and HM + precession separately for the new candidate events; we show the logarithm of the Bayes factors in Table 1. Four of the new candidate events (the ones numbered ) have Bayes factors larger than unity for HM, and the detection of these candidates likely benefited the most from our search containing HM. Most of the remaining events have Bayes factors for HM that are lower than unity, and we judge that the detection of these additional candidates was a result of just an overall improvement of our search’s sensitivity in the high mass/redshift region due to e.g., improved template coverage and background estimation.
We show a few select 2D posteriors of our new candidate events in colored contours in Figs. 1, 2 and 7. We also overplot the contours from events in GWTC-3, 4-OGC and the IAS -only O3 catalogs in light gray. Our search only covers events with Hanford–Livingston coincidence triggers, but we include all binary BH events that are part of the LVK catalogs through O1–O3 in these figures. Note that the posteriors are bi-modal for some events, which show up as two contours in the plot.
We add dashed lines to mark the approximate limits of sensitivity in the top panel of Fig. 1. The dashed line in the top-right of the panel marks the region where the binary starts moving out of band (we take this to happen when the detector frame mass equals , which is the limit of our template bank). The dashed line in the top-left marks the contour of constant detector-frame luminosity for an inspiral-dominated equal-mass binary; it roughly delineates the parameter limits past which binaries are too faint to be observable. It is also worth mentioning that the size of the contours increases for higher masses as the number of cycles in band starts to decrease. In the inspiral phase, the chirp mass is well-measured, but this no longer continues to be the case for high-mass events where the SNR is dominated by the merger and ringdown phase and the total mass of the binary starts to be measured better.
We show the marginalized 1D posteriors for our new candidate events under the aforementioned two different spin priors in Fig. 3. Figs. 10 and 11 in the Appendix show full corner plots from different PE runs. Some of the events in Fig. 3 (e.g., the ones numbered ) have posteriors that show significant support for large, positive effective spin under the uniform in prior. The isotropic spin prior recovers a lower likelihood for these events, since the prior pulls the sampler away from the high solution.
GW191228_195619 has the largest evidence for HM () and also for the combination of HM+precession (). We show a corner plot corresponding to this event in Fig. 4 for our three PE runs, and we indeed see significant support for both HM and precession. Note that for the (green) and + HM (orange) cases, the distribution of the primary component’s in-plane spin component is that of its prior conditioned on the other parameters. For these cases, we set the in-plane spin components to zero when evaluating likelihoods. Within 90% CL, we find to be . Note also that the total SNR (110) recovered by our non-precessing Hanford+Livingston search is significantly smaller than obtained from a PE run which includes both the effects of orbital precession and data from the Virgo detector. Therefore, we conjecture that the significance of this event could be improved in future searches that include the effects of precession and/or Virgo data.
We also compare our results against PE runs using the IMRPhenomXPHM approximant Pratten et al. (2021) (under the two different MSA Chatziioannou et al. (2017) and NNLO Hannam et al. (2014) prescriptions for evolving the spins) in Fig. 9 in the Appendix; we find good agreement in most of the cases. It would be interesting to compare the PE for these events with NR surrogate models (e.g. Islam et al. (2023)) in the future.
We will discuss interesting astrophysical properties of our new candidate events in Section IV below.
III.2 Comparison to the IAS (2, 2)-only catalogs
In Table 2, we report the candidate events with from our pipeline which have already been reported in the previous -only IAS searches in the O3 data Olsen et al. (2022); Mehta et al. (2023). The full comparison catalog including lower candidates is given in Table 6. Overall, we recover 6 events with .
We caution against directly comparing the sensitivities of the -only and + HM pipelines from Table 6 as there are multiple differences: () including HM in the search, () using a different astrophysical prior (see Section II.4) () additional glitch cleaning at high masses (see Section II.3). It is worth mentioning though that we recover the events which were reported in previous IAS catalogs with , except GW190704_104834 (this event was in BBH-0 in O22 and we checked that the difference in the astrophysical priors between the pipelines was large for that bank). We collect a larger number of triggers in the HM search as compared to a -only search, so we raised the collection threshold to mitigate the increased memory requirements. This caused us to miss low-SNR events that were reported by -only pipelines with lower collection thresholds in the previous literature. We further discuss this point in Section III.3.2 below.
One notable event in Table 2 is GW190711_030756. The IFAR of this event in our HM search is 38.7 yr as opposed to yr in the -only search O22 (after accounting for the trials factor in O22 due to the number of banks searched). We show a corner plot for this event in Fig. 5. Indeed, GW190711_030756 also shows a significant evidence for higher modes and our search recovers an additional 10 points in SNR compared to the -only search. See Fig. 3 and section III A.2 of O22 for a detailed discussion of the properties of this event.
III.3 Comparison to the LVK and OGC -only catalogs
In Table 3, we compare our pipeline’s results for the O3 Hanford–Livingston coincident events published in the GWTC-3 catalog333https://gwosc.org/GWTC-3/ Abbott et al. (2021). The GWTC-3 catalog combines results from four pipelines: three that search using templates which only contain the mode: GstLAL, Multi-Band Template Analysis (MBTA) and PyCBC Sachdev et al. (2019); Usman et al. (2016); Andres et al. (2022); and one that searches for transient signals with minimal assumptions about sources, cWB Klimenko et al. (2016). We also include the comparison with the 4-OGC catalog taken from the summary data in the repository444https://github.com/gwastro/4-ogc/blob/master/search/4OGC_top.txt listed in Nitz et al. (2021). Note that the quoted GWTC-3 squared SNR, , is from LVK’s PE runs for the events. The search SNR is generally lower, and varies between the different pipelines that the LVK uses.
We broadly recover the confident LVK detections, except some vetoed/missed events which are discussed in more detail in the subsections below. Overall, 9 events GWTC-3 events in Table 1 were vetoed in our pipeline. 4 GWTC-3 events fell below our trigger collection SNR threshold (GW190917_114630, GW191113_071753, GW200220_061928, GW200308_173609), note that these were confidently detected by only one of the LVK pipelines. Three GWTC-3 events were detected with in our pipeline (GW200208_222617, GW200306_093714, GW200322_091133), note again that these were confidently detected by only one of the LVK pipelines.
At the end of Table 3, we include comparison with the list of 7 events in 4-OGC which are not present in GWTC-3 catalog (from which we recover 4 with and 2 were below our collection SNR threshold). GW190909_114149 is a marginal candidate in GWTC-3 which has from our pipeline. We still include it in Table 3 as it was first a significant candidate in GWTC-2 Abbott et al. (2021), but later updated to be in the marginal list The LIGO Scientific Collaboration et al. (2021). We also checked the overlap of our new candidate event list with the sub-threshold candidate list (i.e., which have but FAR day) released by the LVK collaboration: GWTC-2.1 and GWTC-3 The LIGO Scientific Collaboration et al. (2021); Abbott et al. (2021); we show the corresponding comparison in Table 5 in the Appendix. In the same table, we also show the overlap of our new candidate list with sub-threshold candidates from the IAS -only pipeline Olsen et al. (2022); Mehta et al. (2023).
III.3.1 Vetoed events
Our pipeline includes a series of signal consistency checks in order to veto triggers that are likely to arise from noise transients (glitches) to improve our search sensitivity (see section I of Venumadhav et al. (2019)). The vetoes are designed in a way that they are falsely triggered by real signals with only a low probability (a few percent) if their waveforms are accurately modeled by templates in the bank, and the noise is Gaussian. If a loud signal displays effects which are not present in our templates, like precession or eccentricity, the chances of the veto mis-identifying it as a glitch can be much higher. GW190521_030229, GW191109_010717 and GW200129_065458 were reported to be fairly high SNR with imprints of precession Abbott et al. (2020, 2021), which could be the reason for them being vetoed. For example, the presence of higher modes in GW190521 is disfavored with a Bayes factor of , while a precessing signal is favored with a factor of Abbott et al. (2020).
GW190701_203306 was vetoed due to excess power present after subtracting the best-fit waveform (note that the data around this event was treated for scattered light glitches Abbott et al. (2021)). Out of the veto tests that we apply to events, we noticed GW190828_065509 and GW190930_133541 each failed just one test (namely, the split test with chunk sets, and the test with 20 chunks, respectively).
We noticed a small issue in the pipeline where we artificially veto triggers very close ( s) to the edge of the timeseries files; this led to the event GW190725_174728 being vetoed. We will fix this issue in future searches. The event GW190426_190642 was detected confidently by only one of the LVK search pipelines and got vetoed in our case.
We expect to recover some of loud vetoed events if we also include precession in our template banks. Alternatively, one could formulate a procedure to override vetos or changing their thresholds based on injections with precessing waveforms. We leave exploring these directions to future work.
III.3.2 Other missed/downweighted events
As we discussed earlier in Section II.2, we set a particular SNR threshold () to collect triggers from individual detectors. We do this in order to regulate the number of single-detector triggers that are passed to the more costly coincidence step of the pipeline. If we set the threshold based on the quadrature sum of the SNRs of the (orthogonalized) harmonics, we collect more triggers in a HM search as compared to a -only search with the same threshold. Hence, we raised the collection threshold (see Section II.2.1 for a detailed discussion). Due to this, we missed low-SNR events that were reported by -only pipelines in the previous literature. We label the events which fell below this threshold as “Below ” in the table, where the threshold depends on the number of templates in the bank and is therefore different for different banks. On further investigation, using a smaller value of is feasible with our current search setup, and thus we plan to lower the threshold in future searches.
We do not include the following events in Table 3 because they were detected in Livingston–Virgo or Hanford–Virgo coincidence, or single detector search, all of which we have yet to run: GW190620_030421, GW190630_185205, GW190708_232457, GW190814, GW190910_112807, and GW190925_232845. We also only searched for BBH mergers in this paper, and did not recover the NSBH candidates GW191219_163120 and GW200115_042309 (in addition, the posterior probability for GW200210_092254 peaks in the NSBH region Abbott et al. (2021) and thus it was vetoed in our search). We hope to perform coincident searches which would also include Virgo (and KAGRA) in the future. We also intend to perform a separate search for NSBH mergers including our HM pipeline in a future paper.
GW191222_033537 was detected with in our pipeline but has comparatively a much higher false alarm rate. We checked that it was downweighted due to being in the proximity of a glitch (which led to the event acquiring unphysical amounts of SNR in HM).
IV Interesting astrophysical properties of the new candidate events
In this section, we discuss interesting astrophysical properties of the 14 new candidate events, whose parameters we have reported in Table 1 and in Figs. 1, 2 and 3.
IV.1 Masses in the IMBH range and pair instability mass gap
Black holes roughly within the mass range – are designated as intermediate-mass black holes (IMBHs). Several IMBH candidates have been suggested by electromagnetic observations, but these lack conclusive confirmation Greene et al. (2020). Gravitational waves have the potential to definitively detect objects in this regime, e.g., Abbott et al. (2020); Fragione et al. (2018). In our search, at % credibility, 4 candidate events have total masses in the IMBH range (GW190605_025957, GW190604_103812, GW191228_195619 and GW190530_133833). Among them, GW191228_195619 has the highest total mass (and also has the highest primary mass ), which makes it, by a large margin, the most massive candidate with detected in gravitational wave data so far.
Another interesting regime in BH masses is the pair instability mass gap. This is the regime in which stars undergo pair-instability supernova (where they are completely disrupted and no remnant is left behind). Otherwise, the stars undergo pulsational pair-instability supernova and iteratively lose their mass before collapsing to a black hole. These processes are expected to produce a gap roughly between – in the distribution of BH mass. Note however that the exact boundaries of such a gap are not well determined, as they depend on multiple uncertain factors such as the rate of carbon to oxygen burning in the core , angular momentum transport, metallicity, wind loss, etc. Marchant and Moriya (2020); Farmer et al. (2019); Mehta et al. (2022); Hendriks et al. (2023).
Recent analyses have not found any conclusive evidence for the upper mass gap in the observed BH mass distribution, e.g., Abbott et al. (2023); Roulet et al. (2021). This either challenges the current theoretical models of stellar evolution, or the gap might be filled by black holes formed via alternative mechanisms such as hierarchical mergers Anagnostou et al. (2020); Fragione et al. (2020); Veske et al. (2021); Fragione et al. (2021), or from the gaseous disk channel Tagawa et al. (2020); Samsing et al. (2022). In our search, at 90% CL, 9 candidates have median mass of the primary BH lying between – (# 1, 4, 5, 6, 8, 10, 11, 13, 14 in Table 1). One of our candidates: GW191228_195619 has its secondary BH mass overlapping with the mass-gap range. Our new candidate events could be helpful in estimating the abundance of BH (or rate of BH mergers) in/beyond the pair-instability mass regime. This can also shed light on BH progenitor channels.
IV.2 High redshifts
In our search, we find that 9 of our new candidates have redshift at % CL (# 1, 2, 3, 6, 8, 10, 11, 13, 14 in Table 1). Finding high-redshift BHs can be important to answer questions such as: How does the merger rate of BHs evolve with redshift Fishbach et al. (2018)? Does the merger rate change in a different manner compared to the the star formation rate (which peaks at Madau and Dickinson (2014))? Furthermore, different BH formation channels can have different predictions for the merger rate at high redshift Mapelli et al. (2022); van Son et al. (2022). For example, the chemically homogeneous evolution channel produces BBHs of relatively higher masses and lower redshifts than the common envelope evolution channel Mandel and de Mink (2016). In the dynamical channel, massive BHs can have relatively longer delay times compared to star formation Mapelli et al. (2019). The redshifts of binaries can be also be correlated with other properties of BHs, e.g., spins Biscoveanu et al. (2022); Kushnir et al. (2016); Zaldarriaga et al. (2018). We aim to test some of these models in a future population inference study.
IV.3 Preference for positive effective spins
Measuring spins of the black holes is one of the best ways to distinguish between isolated and dynamical formation channels. In the isolated channel, accretion can lead to spins of BHs which are aligned in the direction of the binary angular momentum (leading to positive values of the effective spin parameter ). Dynamical channels, on the other hand, predict the astrophysical distribution to be symmetric around 0. This is because the the spins BH captured into a binary are not expected to be preferentially aligned along any particular direction. From the PE runs under our fiducial prior, 6 candidates have effective spin at credibility (# 2, 4, 6, 8, 9, 11 in Table 1). Note however that posteriors are significantly dependent on the prior as seen in the comparison in Fig. 3. Though, even under the isotropic spin prior, some of the new candidate events still have mild preference for .
Note that the preference for in our detection catalog does not immediately mean that astrophysics prefers this scenario, as this could be caused by selection effects. One of the selection effects is that positive binaries are louder than their negative counterparts; due to the orbital hangup effect, the positive binaries merge later and have power at higher frequencies Ng et al. (2018); Roulet and Zaldarriaga (2019). This effect is much more prominent for high mass binaries (where a significant portion of signal is masked by the low frequency end of the PSD) as we show in Fig. 6. There is another effect which further downweights high-mass negative signals. Their waveforms have a short duration in band and therefore can be more effectively mimicked by glitches. This degrades the sensitivity to such systems in our search Wadekar et al. (2023b) when we perform the background dependent reweighting in our ranking statistic (see Section II.3.3 and Ref. Wadekar et al. (2023b)). To robustly determine if the true astrophysical distribution also favors at high masses, we plan to a do a population study (where we will use injections with both positive and negative values to determine our pipeline selection function).
IV.4 Asymmetric mass ratio systems
Five of our new candidates have median mass ratio (# 4, 6, 7, 10, 11 in Table 1). Finding sources with asymmetric mass ratios can shed light on the formation channels Abbott et al. (2020b). The globular cluster (dynamical) channel predicts that most merging BBHs have nearly equal masses Rodriguez et al. (2016). This is because higher mass BHs in a globular cluster sink to their cores due to dynamical friction and vice versa for lighter BHs. The caveat is that if clusters can retain merger products, they can produce “second-generation” mergers whose mass ratios can deviate significantly from unity Rodriguez et al. (2019). In the isolated binary channel, in certain cases, an initially asymmetric mass ratio system can evolve toward a more symmetric configuration due to mass accretion episodes Ivanova et al. (2013); Postnov and Yungelson (2014); note that the reverse scenario has been proposed as well (see e.g., Ref. Broekgaarden et al. (2022)). Another interesting property of asymmetric mass ratio and high total mass systems is that subdominant ringdown modes have a higher chance of being observable in such systems, see e.g. Capano et al. (2021).
V Discussion
V.1 Dependence on the choice of astrophysical prior
Our estimate of and our overall IFAR depend on the astrophysical prior used in the search, and that dependence is strongest in the case of near-threshold detections like some of the new candidates in our catalog list in Table 1.
To facilitate further analysis and comparisons between different astrophysical priors, we make our posterior samples and other relevant information public at https://github.com/JayWadekar/GW_higher_harmonics_search. These can be used to reweight our results with other choices of priors (see e.g., Payne et al. (2019)). The biggest difference in our astrophysical prior choice compared to other catalogs Abbott et al. (2021); Nitz et al. (2021) is for the BH spins. We use a prior that is flat in the effective spin parameter . Oue motivation lies in the fact that is often the best-measured spin parameter, and hence we want to give equal weights to all physically possible values of that parameter Zackay et al. (2019). On the other hand, using an isotropic spin prior will significantly down-weight large (positive or negative) effective spins and can make it harder to reweight the results to a different prior. We also provide PE samples for posteriors sampled under the isotropic spin prior in the above repository. Another difference in our priors is for the case of total mass and mass ratios.
It is also worth discussing how our other priors compare with the results of Abbott et al. (2023). We used a search prior uniform in , whereas, Ref. Abbott et al. (2023) modeled the distribution as a power law and found . We used a simple power law prior for the detector frame total masses and did not include any separate prior over redshift. Ref. Abbott et al. (2023) modeled the source primary mass distribution as a power law with supplemented by a Gaussian peak at . Ref. Abbott et al. (2023) modeled the merger rate as and found . The motivation for our prior choices was again to make it easy to reweight our results (both and overall IFAR values) later in a hierarchical population analysis (especially for the high mass and low mass ratio parameter space), see e.g., Roulet et al. (2020).
A full population analysis not only requires the PE samples but also the of each event Roulet et al. (2020). For this reason, we also provide a file, IAS_HM_O3_triggers.hdf, in the repository that contains the full list of triggers (including the sub-threshold ones). Just for reference, we report subthreshold events () in Table 7 in the Appendix.
V.2 What is the gain/loss in sensitivity due to including HM in the search?
Unfortunately, we cannot currently answer the above question by comparing the results from our search to those from the previous IAS -only searches Olsen et al. (2022); Mehta et al. (2023). This is because our search not only includes the effect of HM, but also includes the improvements (for all modes) due to downweighting noise transients (see Section II.3 and Wadekar et al. (2023b)). Furthermore, compared to earlier, we use a different astrophysical prior (see section II.4) and use a different waveform approximant for modeling the waveforms (IMRPhenomXAS Pratten et al. (2020) instead of IMRPhenomD Khan et al. (2016)). For us to isolate the effect of HM, we will need to rerun our search (ideally on injections) with all the settings to be the same except having only modes in the template banks. We leave this study to a future work.
Previous studies have noted that the larger variety in waveform morphologies (brought by the inclusion of higher modes) can lead to a loss of sensitivity in the search, as the number of background triggers is increased, see e.g., Capano et al. (2014); Chandra et al. (2022). This can be understood from the fact that, with higher modes, the shape of the waveform also depends on the inclination and orbital phase, and these additional degrees of freedom increase the trials factor when finding the best-matching template. In our search, we do not blankly add additional degrees of freedom to the templates throughout the parameter space. Instead, our new detection statistic (see Section II.2 and Wadekar et al. (2023b)) includes HM while simultaneously marginalizing the likelihood over the extrinsic parameters (which are the same as in the -only search). This marginalization naturally downweights fine-tuned solutions, approaching the optimal statistic and preventing an undue increase of the trials factor (e.g., the edge-on solutions get naturally downweighted when we marginalize over distance, as their observable volume is smaller).
In our ranking statistic for the triggers, we also marginalize over the ratios of the SNR amplitudes of HM: and . In this integral, we use a prior distribution for derived from our astrophysical prior. Note that the prior on can be very different for templates in different regions of parameter space. For example, the median of our prior for is 0.04 for Bank 0 (a low-mass bank), while it is 0.3 for Bank 14 (a high-mass bank). If we did not include such a prior, we would indeed expect to encounter a large loss of sensitivity on including HM. Comparing the performance of our search on events detected in previous -only catalogs (see Tables 3 and 6), we can confirm that our search sensitivity for these events has not been significantly degraded and is roughly comparable.
We also find that in our new search, the significance of a few previously reported marginal events is increased significantly (GW190711_030756, GW200216_220804). Especially for the case of GW190711_030756, we suspect that the increase in sensitivity is due to addition of HM, but to test this we plan to do an injection study in the future.
The downside of using a prior distribution on is that it is dependent on our choice of astrophysical prior (which is currently not precisely known). Especially, the mass ratio distribution is not very well constrained and affects our prior for . Using a different distribution for than the true astrophysical distribution can indeed lead to some loss of sensitivity. We chose a broad prior (flat in ) in our search, with the motivation that such a prior can be easily reweighted later in a hierarchical population analysis.
Even though including HM can cause some loss in sensitivity for near-threshold equal mass systems, note that detecting even a few events with asymmetric-mass ratios can inform our population estimates much more than detecting a slightly larger number of equal-mass systems. It is worth drawing an analogy here to the case of spin distribution of templates used in -only searches. Neglecting templates with negative indeed improves the search sensitivity to near-threshold systems, but finding even a few systems can greatly help us distinguish between progenitor channels of BHs. We thus argue the goal of a search should not solely be to maximize the total number of detections, but to holistically cover most of the interesting regions of the binary parameter space. A similar argument also holds for a precessing search, where detecting even a few events with significantly misaligned spins can shed more light on the formation channels of binaries as compared to detecting a larger number of near-threshold aligned/zero spin mergers.
VI Conclusions
Nearly all gravitational searches performed till now only include the quadrupole mode of the full GW waveform predicted by general relativity. In our companion papers Wadekar et al. (2023a, b), we develop new efficient methods to introduce higher-order modes (HM) in our search pipeline (section II). This helped us perform the first search over the full binary black hole parameter space (i.e., over all binary inclinations) with including HM in our templates. Compared to our previous pipeline, we introduced additional improvements to remove/downweight noise transients (especially at low frequencies), which further improved our sensitivity to high mass and high redshift systems (section II.3).
We searched over public data from the third LIGO–Virgo observing run and found 14 new candidate events with (see Table 1 and section III). Some of the new candidate events have several interesting astrophysical properties (section IV) such as: () masses in the IMBH range and pair instability mass gap, () high redshifts, () preference for positive effective spins, () asymmetric mass ratios.
By simply summing over the complements of values of our new candidate events from Table 1, we expect that roughly 4 of the 14 new events are noise transients instead of being astrophysical signals. However, we note that estimates of and source parameters depend on the choice of prior, and we discuss this dependence in section V555We provide our posterior samples at https://github.com/JayWadekar/GW_higher_harmonics_search, which can be used to reweight our results for a different astrophysical prior..
We also compared our pipeline’s results for the events published in previous IAS catalogs (Tables 2 & 6), and in the GWTC-3 and 4-OGC catalogs (Table 3). We broadly recovered the high-significance events, except some events which were either vetoed by our pipeline or fell below our trigger collection SNR threshold, which we discuss in detail in section III.3 (we also mention some possible future improvements in this regard).
Our methodology for efficiently including realistic effects like HM in GW templates will continue to be useful for high-precision searches in upcoming and future gravitational wave experiments Fairhurst et al. (2023); Maggiore et al. (2020); Evans et al. (2023); Kawamura et al. (2020); Amaro-Seoane et al. (2017); Akutsu et al. (2021). One particular example in which HM will be especially useful is when searching for extreme-mass-ratio inspirals in LISA Zhang et al. (2023); Baibhav et al. (2020).
VI.1 Future work
We plan to inject simulated gravitational-wave signals in the data and compare the fraction of injections recovered by our pipeline. The injections will help us determine the gain/loss in volume-time sensitivity in different regions of the parameter space due to including HM in our search. The injections will also help us gauge the level of improvement due to the additional glitch treatment methods that we described in section II.3. We plan to use the injection parameters from Abbott et al. (2021), which will help us perform an apples-to-apples comparison of our sensitivity with other pipelines. Having injections will also help quantify the selection function of our current pipeline and help in population inference studies. The new events reported here have moderate IFARs (the best event has yr), thus a methodology to combine confident and marginal events in population inference will be required Roulet et al. (2020). We expect our sensitivity to be higher than other searches for asymmetric mass ratio events due to addition of HM. Our sensitivity could also be higher for high-mass and high-redshift regions due to using improved glitch treatment methods and adding HM. Note that, in certain regions of the parameter space, having higher sensitivity can improve population estimates even if we have null/marginal detections. We note some interesting population inference questions in section IV.
Currently, nearly all searches also ignore the effects of precession or eccentricity in their waveforms. Some of the challenges for including these effects are similar to the ones for HM: significant increase in the template bank size and loss of sensitivity due to collecting additional background. We hope to develop methods similar to the ones implemented in our search to help tackle those searches in the future.
Acknowledgements.
We thank Horng Sheng Chia, Ajith Parameswaran, Will Farr, Vishal Baibhav, Katerina Chatziioannou, Vicky Kalogera, Sylvia Biscoveanu, Amanda Farah, Daniel Holz, Salvatore Vitale, Tom Edwards, Zoehyr Doctor, Suvodip Mukherjee and Mukesh Singh for helpful discussions. DW gratefully acknowledges support from the Friends of the Institute for Advanced Study Membership and the Keck foundation. TV acknowledges support from NSF grants 2012086 and 2309360, the Alfred P. Sloan Foundation through grant number FG-2023-20470, the BSF through award number 2022136, and the Hellman Family Faculty Fellowship. MZ is supported by NSF 2209991 and NSF-BSF 2207583. BZ is supported by the Israel Science Foundation, NSF-BSF and by a research grant from the Willner Family Leadership Institute for the Weizmann Institute of Science. This research was also supported in part by the National Science Foundation under Grant No. NSF PHY-1748958. We also thank ICTS-TIFR for their hospitality during the completion of a part of this work. This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain.Appendix A Additional plots and tables
Sr. | New candidate | This search | GWTC-3 sub-threshold | IAS 22-only sub-threshold | ||||||
No. | event | IFAR | IFAR | |||||||
1 | GW190605_025957 | 75.7 | ||||||||
2 | GW190806_033721 | 58.4 | ||||||||
3 | GW190524_134109 | 60.0 | ||||||||
4 | GW191113_103541 | 50.8 | Veto | |||||||
5 | GW190615_030234 | 65.9 | ||||||||
6 | GW190604_103812 | 46.8 | ||||||||
7 | GW191228_195619 | 98.9 | Veto | |||||||
8 | GW200304_172806 | 52.5 | ||||||||
9 | GW190530_030659 | 60.0 | ||||||||
10 | GW190511_163209 | 46.9 | ||||||||
11 | GW200210_100022 | 53.7 | ||||||||
12 | GW200301_211019 | 60.9 | ||||||||
13 | GW190911_195101 | 65.7 | ||||||||
14 | GW190530_133833 | 53.3 |
(O3a) Name | Bank | Best fit template | IFAR (yr) | ||||||
GW190711_030756 | 7,1 | ||||||||
GW190707_083226 | 7,1 | ||||||||
GW190818_232544 | 8,1 | ||||||||
GW190906_054335 | 7,1 | ||||||||
GW190814_192009 | 11,0 | ||||||||
GW190821_124821 | 0,1 | ||||||||
GW190704_104834 | 0,0 | ||||||||
GW190920_113516 | 0,0 | ||||||||
GW190910_012619 | 5,0 | ||||||||
(O3b) Name | Bank | IFAR (yr) | |||||||
GW200109_195634 | 10,0 | ||||||||
GW191228_085854 | 0,2 | ||||||||
GW200316_235947 | 0,0 |
Candidate | Bank | Best fit template | IFAR (yr) | |||||||
per bank | overall | |||||||||
GW190708_211916 | 8,1 | |||||||||
GW190907_111633 | 12,0 | |||||||||
GW191121_042622 | 13,0 | |||||||||
GW190810_180039 | 6,2 | |||||||||
GW200304_182240 | 11,0 | |||||||||
GW200104_184028 | 8,1 | |||||||||
GW191127_114537 | 13,0 | |||||||||
GW190423_083253 | 10,0 | |||||||||
GW200324_083227 | 7,1 | |||||||||
GW200312_093903 | 11,0 | |||||||||
GW200308_231646 | 8,1 | |||||||||
GW190620_221103 | 7,1 | |||||||||
GW200211_113417 | 11,0 | |||||||||
GW200128_115507 | 8,1 | |||||||||
GW190606_061230 | 0,2 | |||||||||
GW190928_163629 | 4,2 | |||||||||
GW200208_194804 | 13,0 | |||||||||
GW200217_211137 | 9,0 | |||||||||
GW190426_180646 | 7,1 | |||||||||
GW200307_084747 | 9,0 | |||||||||
GW200216_200717 | 8,1 | |||||||||
GW190403_061805 | 6,0 | |||||||||
GW200226_083409 | 6,2 | |||||||||
GW190521_071322 | 7,1 | |||||||||
GW190908_161547 | 7,1 | |||||||||
GW190911_021724 | 4,2 | |||||||||
GW200303_074125 | 7,1 | |||||||||
GW190526_231941 | 7,1 | |||||||||
GW200319_022728 | 8,1 | |||||||||
GW200303_083614 | 0,1 | |||||||||
GW191206_163102 | 13,0 | |||||||||
GW191203_034221 | 3,1 | |||||||||
GW200305_225311 | 3,1 | |||||||||
GW200126_225208 | 7,1 | |||||||||
GW191209_143052 | 0,1 | |||||||||
GW200128_113935 | 4,2 | |||||||||
GW200201_030733 | 4,2 | |||||||||
GW200202_160037 | 4,2 | |||||||||
GW191209_191159 | 14,0 | |||||||||
GW200212_101822 | 0,3 | |||||||||
GW190911_025827 | 14,0 | |||||||||
GW200303_013439 | 0,2 | |||||||||
GW190802_095334 | 5,2 | |||||||||
GW191203_121111 | 14,0 | |||||||||
GW190928_223455 | 2,0 | |||||||||
GW191227_044656 | 14,0 |

### v2

@@
 Nearly all of the previous gravitational wave (GW) searches in the LIGO–Virgo data included GW waveforms with only the dominant quadrupole harmonic, i.e., omitting higher-order harmonics which are predicted by general relativity.
-Based on the techniques developed in Wadekar et al.
-(2023a, b), we improve the IAS pipeline by () introducing higher harmonics in the GW templates, () downweighting noise transients (“glitches”) to improve the search sensitivity to high-mass and high-redshift binary black hole (BBH) mergers.
-We find 14 new candidate BBH mergers with on running our pipeline over the public LIGO–Virgo data from the O3 run (we use the detection threshold as the astrophysical probability, , being over 0.5, following the approach of other pipelines).
-We also broadly recover the high-significance events from earlier catalogs, except some which were either vetoed or fell below our SNR threshold for trigger collection.
+We improved the IAS pipeline by efficiently introducing higher harmonics in the GW templates using the techniques in Wadekar et al.
+(2023, 2024).
+Using the IAS-HM111The pipeline is publicly available at https://github.com/JayWadekar/gwIAS-HM pipeline on the public LIGO–Virgo data from the O3 run, we find 11 new candidate BBH mergers with (we use the detection threshold as the astrophysical probability, , being over 0.5, following the approach of other pipelines).
+We broadly recover the high-significance events from earlier catalogs, except a few which were vetoed.
+We also find that including higher harmonics in our search raises the significance of a few previously reported marginal events (e.g., GW190711_030756).
 […]
@@
 (2021); Abbott et al.
-(2021) and by other independent pipelines Venumadhav et al.
+(2021, 2022); Kumar and Dent (2024) and by other independent pipelines Venumadhav et al.
 (2019, 2020); Olsen et al.
@@
 (2019, 2020, 2021, 2021); Chia et al.
+(2024); Mehta et al.
+(2023); Koloniari et al.
 (2024) who access the data made publicly available through the Gravitational Wave Open Science Center (GWOSC) Vallisneri et al.
@@
 Approximate formulae for the relative strength of frequency-domain amplitudes of two of the higher-order modes (HM) are given by Wadekar et al.
-(2023a)
+(2023)
 where is the mass ratio (), is the orbital frequency and is the binary inclination ( corresponds to the edge-on configuration).
-These harmonics have a number of important applications for characterizing properties of BBHs.
+These higher harmonics have a number of important applications for characterizing properties of BBHs.
 If one uses the mode waveform alone in inferring the BBH properties, we have the following important degeneracies between different binary parameters: luminosity distance and ; and effective spin (); polarization and initial orbital phase ().
@@
 (2014).
-This follows from the amplitudes of HM having a different dependence on and as compared to that of the mode (see Eq.
-(1)), and similarly the phases of HM have different dependence on .
+This follows from the amplitudes of HM in Eq.
+(1) having a different dependence on and as compared to that of the mode, and similarly the phases of HM have different dependence on .
 It is worth mentioning that the degeneracy between distance and inclination is currently one of the most important limiting factors in inferring the Hubble parameter from GW systems and HM can be crucial in breaking this degeneracy.
@@
 (2020); Capano et al.
-(2021), and to study properties of the merger remnant such as the recoil kick.
-Multiple waveform models including HM have been recently made available London et al.
-(2018); García-Quirós et al.
+(2021); Islam et al.
+(2021a, b), and to study properties of the merger remnant such as the recoil kick Calderón Bustillo et al.
+(2018).
+Multiple waveform models including HM have been recently made available (e.g., London et al.
+(2018); Garc´ıa-Quirós et al.
 (2020); Mehta et al.
-(2017), and they are currently being used ubiquitously in parameter estimation studies.
-However, HM are still not included in current search pipelines (as of now, the only exception is Ref.
-Chandra et al.
-(2022), which searched for binaries with nearly edge-on inclinations).
+(2017); Ramos-Buades et al.
+(2023)), and they are currently being used ubiquitously in parameter estimation studies.
+However, HM are still not included in current search pipelines (as of now, the only exception is Chandra et al.
+(2022), which however only searched for binaries with nearly edge-on inclinations).
 The reason behind this is that for making template banks with the mode, one only needs to sample over the masses and spins of black holes ().
@@
 To tackle the first issue, we found more efficient ways of including HM in our template banks (which we presented in detail in our first companion paper Wadekar et al.
-(2023a)).
+(2023)).
 Regarding the second issue, we devise new detection statistics which includes marginalization over the amplitudes and phases of higher modes in our second companion paper Wadekar et al.
-(2023b).
+(2024).
 We present an overview of these new techniques in Section II.
@@
 Moreover, due to the shape of the current detector power spectral density (PSD), the signal-to-noise ratio (SNR) contribution of HM is additionally boosted at high masses and redshifts.
-In this regime, the mode can cut off below the detector’s sensitive band (the detector PSD sharply increases below Hz) and HM, being at higher frequencies, naturally have an increased contribution to the SNR.
+In this regime, the lower-frequency mode can cut off below the detector’s sensitive band (the detector PSD sharply increases below Hz) and HM (being at higher frequencies) naturally have an increased contribution to the SNR.
 The background corresponding to the high-mass regime is however heavily dominated by short-duration noise transients (called instrumental “glitches”).
@@
 We therefore introduce additional methods in the IAS pipeline to downweight the background at high masses.
-These methods are detailed in our second companion paper Wadekar et al.
-(2023b) and we present an overview in Section II.3.
+Some of these methods are detailed in our second companion paper Wadekar et al.
+(2024) (hereafter W24) and we present an overview in Section II.3.
 Running our new search pipeline over the public LIGO–Virgo data from the third observing run (O3), we found new candidate events with properties shown in Figs.
@@
 II Pipeline improvements
-Here, we present an outline of the additions that we made to the -only IAS pipeline Olsen et al. (2022); Venumadhav et al. (2019).
+Here, we present an outline of the additions that we made to the previous -only version of the IAS pipeline Olsen et al. (2022); Venumadhav et al. (2019).
 II.1 Adding HM in template banks
 In this sub-section, we summarize the methodology for constructing template banks for HM from our companion paper: Wadekar et al.
-(2023a).
+(2023).
 For each mode template, we use a combination of post-Newtonian formulae and machine learning tools to model normalized aligned-spin templates corresponding to the and modes.
 We then filter the data separately with the three templates and store the complex SNR timeseries , see figure 1 of Wadekar et al.
-(2023a).
+(2023).
 As a result, our HM search matched-filtering cost is just that of a -only search (in contrast to being as in previous HM search methods, where the different modes are combined and templates sample over inclination and initial orbital phase Chandra et al.
@@
 Note that the division into banks and subbanks is only to facilitate covering the parameter space with templates of the form that we use, and the final ranking statistic we assign to each trigger is weighed against those of all other triggers (over the entire parameter space) to estimate its global significance.
-Other than adding HM, we also improved our template banks by using a machine learning tool called random forest regressor (which helps us reduce the dimensionality of our template banks and enables us to search with 3040% fewer templates than before) Wadekar et al.
-(2023a).
+Other than adding HM, we also improved our template banks by using a machine learning tool called random forest regressor (which helps us reduce the dimensionality of our template banks and enables us to search with 3040% fewer templates than before), see Fig.
+5 of Wadekar et al.
+(2023).
 II.2 New statistics for combining SNR timeseries from different harmonics
-We present the techniques for combining the timeseries in our companion paper (Wadekar et al.
-(2023b)) and here we provide a brief overview.
+We present the techniques for combining the timeseries in our companion paper (W24) and here we provide a brief overview.
 Combining SNR timeseries of different modes incorrectly can lead to a large loss of sensitivity of the pipeline Capano et al.
 (2014).
-We use the Neyman–Pearson lemma to develop an optimal statistic to combine the SNR timeseries for different harmonics; our detection statistic is the ratio of the Bayesian evidence under the signal and the noise hypothesis.
+We use the Neyman–Pearson lemma to develop an optimal statistic to combine the SNR timeseries for different harmonics; our detection statistic is the ratio of the Bayesian evidence under the signal and the noise hypothesis Neyman and Pearson (1933).
 Ideally, one would marginalize over all the intrinsic and extrinsic binary parameters to calculate the evidence under the signal hypothesis.
 We develop two different approximate versions of the statistic to be used in different parts of our pipeline as detailed in the sub-sections below.
-II.2.1 Single-detector marginalized statistic
+II.2.1 Single-detector collection statistic
 We first developed an approximate statistic to collect individual detector triggers.
@@
 For collecting the triggers in a -only search, one can simply use .
-For the HM case, if we naively use , the number of collected triggers increases by a factor of over (using the same value of as the -only case Wadekar et al.
-(2023b)).
-To ameliorate this, we develop in Ref.
-Wadekar et al.
-(2023b) a new marginalized statistic () which exploits the fact that the values of and are bounded for physical signals (as opposed to Gaussian noise triggers where the values and can be arbitrarily large).
-The statistic includes marginalization over physically possible amplitude ratios of higher harmonics.
-To make the computation of faster, we maximize (instead of marginalizing) over the phases of HM and also over all the remaining intrinsic and extrinsic binary parameters (although this can be improved in future studies).
-We collect single-detector triggers using and we use a larger value of the collection threshold as compared to the -only case.
-The downside of increasing the value of is that low-SNR events can fall below our threshold bar.
-This is indeed the reason behind our pipeline missing some of the low-SNR events reported by -only pipelines in the previous literature, see Section III.3 below for more details.
-The particular choice of that we made in the current paper is currently not a bottleneck for our search and can be improved in the future.
+For the HM case, if we naively use , the number of collected triggers increases by a factor of over (using the same value of as the -only case (W24)).
+To ameliorate this, we develop a new marginalized statistic to collect triggers ( in W24).
+We exploit the fact that the values of and are bounded for physical signals (as opposed to Gaussian noise triggers where the values and can be arbitrarily large), see Fig.
+2 of (W24).
+We therefore marginalize over the space of physically possible amplitude ratios of higher harmonics using a pre-computed prior samples for (which are derived from our astrophysical prior).
+Note that the prior on can be very different for templates in different regions of parameter space.
+For example, the median of our prior for is 0.04 for Bank 0 (a low-mass bank), while it is 0.3 for Bank 14 (a high-mass bank).
+If we did not include such a prior, we would indeed expect to encounter a large loss of sensitivity on including HM.
+To make the computation of faster, we maximize (instead of marginalizing) over the phases of HM and also over all the remaining intrinsic and extrinsic binary parameters.
 II.2.2 Coherent marginalized statistic
@@
 Our coherent ranking statistic thus includes marginalization over extrinsic parameters in a similar way to the IAS -only search methodology given in Appendix D of Olsen et al.
-(2022) (hereafter O22) except generalized to account for higher modes.
+(2022) (hereafter O22), except generalized to account for higher modes (see Roulet et al.
+(2024); Leslie et al.
+(2021) for details).
 In addition, we also include marginalization over the amplitude of HM (which effectively corresponds to marginalizing over mass ratios); the only parameters that we currently do not marginalize over are the intrinsic binary parameters corresponding to the 22 waveform, as we use only the best-fit 22 template in our calculations.
-Details regarding our coherent statistic will be given in upcoming papers Wadekar et al.
-(2023b); Roulet et al.
+Details regarding our coherent statistic are provided in our companion papers Wadekar et al.
+(2024); Roulet et al.
 (2024).
 II.3 Removing/downweighting noise transients
-In addition to the changes involving higher modes, we also improve the sensitivity of our search for high masses compared to previous IAS pipelines by developing new tools.
-We also discuss in this sub-section methods to tailor the background reduction techniques in our -only pipeline for the case of HM.
-II.3.1 Band eraser tool for data preprocessing
-In our -only pipeline, in a data pre-processing step (before we matched-filter with templates), we identify regions of bad data segments which have excess power localized to particular frequency ranges and timescales (see section C of Venumadhav et al.
-(2019)).
-We then remove the data in the flagged regions (i.e., make “holes”) and then fill the holes using an “inpainting filter” (see figure 6 of Zackay et al.
+In addition to the changes involving HMs, we also improve the sensitivity of our search for high masses compared to previous IAS pipelines by developing new noise reduction tools.
+II.3.1 Trigger Inference using Extended strain representation (TIER)
+We developed a machine learning (ML) framework called TIER (Trigger Inference using Extended strain Representation) for improving the sensitivity of any generic pipeline Wadekar et al.
+(a).
+Typically, search pipelines only use a small region of strain data in the vicinity of a candidate signal to construct the detection statistic.
+However, extended strain data ( s) in the candidate’s vicinity can also carry valuable complementary information.
+We included this extended information into our detection statistic using ML classifiers (see Fig.
+1 of Wadekar et al.
+(a)).
+Using TIER, we found up to 20% improvement in sensitive volume of our pipeline, with improvements being concentrated in regions of high masses and unequal mass ratios.
+We refer the reader to Wadekar et al.
+(a) for further details regarding TIER.
+II.3.2 Band eraser tool for data preprocessing
+In our -only pipeline, in a data pre-processing step (before we matched-filter with templates), we first identify regions of noisy data segments (see section C of Venumadhav et al.
+(2019)), and remove the data in the flagged regions (i.e., make “holes”).
+We then fill the holes using an “inpainting filter” (see Fig.
+6 of Zackay et al.
 (2021)).
@@
 Sometimes, however, the noise transients are only localized to particular frequency bands (particularly the low-frequency end 100 Hz) and do not affect astrophysical signals which dominantly have power in a different frequency band.
-Therefore, before making holes in entire time-chunks, we try to smoothly remove bad time-frequency regions (or bands) in the spectrogram.
-For this, we develop a new tool we call band eraser.
-We divide the spectrogram into multiple bands (each with dimensions ).
+Therefore, before making holes in entire time-chunks, we try to smoothly remove bad time-frequency regions (or bands) in the spectrogram by developing a new tool called band eraser.
+First, we divide the spectrogram into multiple bands (each with dimensions ).
 We further split each band into multiple chunks (of ) and measure the number of chunks which have power above a certain threshold.
-If this number is above a certain threshold, we remove the entire band from our analysis.
-We calculate this threshold using Poisson statistics, making sure that, on average, in Gaussian noise, one band is removed per s (we also add an additional safety factor to the threshold to avoid removing gravitational wave signals).
-We also noticed that some of the noise transients had power spread over long timescales ( s) and thus were sometimes missed by the localized excess-power tests (with timescales s) that we had used earlier to identify regions for making holes (see Table II of Venumadhav et al.
-(2019)).
-Empirically, we noticed that such disturbances predominantly pollute the low-frequency ( Hz) end, and therefore particularly reduced the sensitivity of our pipeline to high-mass mergers.
-We find that using the band eraser also helped in mitigating some such disturbances.
-We leave further discussion of the band eraser tool to Ref.
-Wadekar et al.
-(2023b).
-II.3.2 Veto procedure
+If this number of chunks in a given band is above a certain threshold, we remove the entire band from our analysis.
+We leave further discussion of the band eraser tool to W24.
+II.3.3 Improving signal consistency vetos using ML
 The -only IAS pipeline included a series of signal consistency checks in order to veto triggers that are likely to arise from noise transients, in order to improve the search sensitivity (see section I of Venumadhav et al.
@@
 We aim to modify these tests to also check the consistency of the and parts of the signals in the future.
+One of the downsides of our veto tests is that if a loud signal displays effects which are not currently present in our templates (e.g.
+precession or eccentricity), the chances of a veto test mis-identifying it as a glitch are much higher (e.g.
+GW190521 was a high-mass merger which got vetoed in a previous IAS search O22 as it shows a mild evidence for precession).
+To remedy this, we run our pipeline on precessing GW signals injected in the data.
+We then use a ML classifier to distinguish between properties of GW signals which got vetoed, as opposed to the properties of the noise transients which got vetoed.
+The ML classifier therefore identifies the precise combination of signal consistency tests which are more likely to fire when a GW signal is precessing.
+The ML classifier output probability is then used to set a criteria for overriding the vetos in a select few cases.
+This helps us recover loud precessing signals and increases the volume-time sensitivity of our pipeline by 10% across nearly all of the parameter space.
+We leave further discussion of the ML-based veto overriding procedure to a future paper Wadekar et al.
+(a).
 In previous quadrupole-only searches, we reduced the noise background by accounting for the non-stationarity of the PSD Zackay et al.
 (2021).
-In the procedure, we match-filtered a reference template from the bank with the data and registered its SNR .
-We then estimated the variation in PSD as a function of time by averaging the power within rolling windows of length s.
-In this search, we calculate the correction for the PSD drift in the same way as before by only using waveforms from our template bank (and not including HM waveforms).
-In the limit in which the fluctuations of the PSD are described by a scalar over all frequencies, this procedure returns an unbiased estimate of the drift correction.
-II.3.3 Ranking statistic
-Even after performing the above steps, there are still some noise triggers which survive.
-We use the ranking statistic as a final tool to downweight such triggers.
+In this procedure, we matched-filter a reference template from the bank with the data and register its SNR .
+We then estimate the variation in PSD as a function of time by averaging the power within rolling windows of length s.
+In the limit in which the fluctuations of the PSD over all frequencies are described by a single scalar, this procedure returns an unbiased estimate of the drift of the PSD.
+In our HM search, we calculate the correction for the PSD drift in the same way as before by only using a reference waveform from the template bank for calculating .
+We currently do not include (3,3) or (4,4) waveforms in estimating PSD drift and leave this direction to future work.
+II.3.4 Ranking statistic
+Even after performing the above noise-removal steps, there are still some noise triggers which survive.
+Our ranking statistic is the final tool we can use to downweight such triggers.
 Beyond upweighting triggers which have good coherence among multiple detectors, an important job of the ranking statistic is to also downweight triggers based on the probability with which similar triggers occur in the background.
@@
 However, we empirically measure by constructing a histogram of the values of background triggers associated with the template .
-In our previous papers, we labelled this term as the “rank function” (), see section J of Ref.
-Venumadhav et al.
-(2019).
 In a similar way for the HM case, if the noise were Gaussian, separately follow distributions (note that we orthogonalize templates of different harmonics), and their sum thus follows a distribution with 6 d.o.f.
-This motivated us to make the approximation and again measure this probability distribution empirically by constructing a 1D histogram of values of background triggers associated with the template .
-We plan to explore beyond the current assumption in a future study.
+This motivated us to make the approximation and again measure this probability distribution empirically by constructing a 1D histogram of values of background triggers associated with the template (see W24).
 Note that in the previous IAS studies, is assumed to be the same for all templates in a subbank.
@@
 (2017); Wadekar et al.
-(2023b).
-In order to optimize our analysis, we follow the technique in Ref.
-Wadekar et al.
-(2023b), where we first cluster the templates within each sub-bank based on their sensitivity to glitches (as quantified empirically by the relative fraction of loud and faint triggers belonging to each template).
+(2024).
+In order to optimize our analysis, we follow the technique in W24, where we first cluster the templates within each sub-bank based on their sensitivity to glitches (as quantified empirically by the relative fraction of loud and faint triggers belonging to each template).
 We then separately construct the histograms of background triggers for each template group to estimate the noise probabilities of triggers associated to templates within that group.
-In a future paper, we plan to quantify the improvement in volume-time sensitivity due to the background reduction procedures outlined in this section, by injecting simulated gravitational-wave signals in the data and comparing the fraction of injections recovered by our pipeline.
 | Sr. | New candidate | Bank | PE (90% credibility) | (evidence) | IFAR (yr) | |||||||||
 | No. | event | HM | HM+Prec. | per bank | overall | |||||||||
-1 | GW190605_025957 | 13,0 | ||||||||||||
-2 | GW190806_033721 | 8,1 | ||||||||||||
-3 | GW190524_134109 | 10,0 | ||||||||||||
-4 | GW191113_103541 | 9,0 | ||||||||||||
-5 | GW190615_030234 | 10,0 | ||||||||||||
-6 | GW190604_103812 | 12,0 | ||||||||||||
-7 | GW191228_195619 | 14,0 | ||||||||||||
-8 | GW200304_172806 | 8,1 | ||||||||||||
-9 | GW190530_030659 | 4,2 | ||||||||||||
-10 | GW190511_163209 | 8,0 | ||||||||||||
-11 | GW200210_100022 | 9,0 | ||||||||||||
-12 | GW200301_211019 | 1,2 | ||||||||||||
-13 | GW190911_195101 | 12,0 | ||||||||||||
-14 | GW190530_133833 | 12,0 |
-Event | Bank | PE (90% credibility) | (evidence) | IFAR (yr) | |||||||||
-HM | HM+Prec. | per bank | overall | ||||||||||
-GW190711_030756 | 7,1 | ||||||||||||
-GW190707_083226 | 7,1 | ||||||||||||
-GW200109_195634 | 10,0 | ||||||||||||
-GW190818_232544 | 8,1 | ||||||||||||
-GW190906_054335 | 7,1 |
-GWTC-3 events | Bank | IFAR (yr) | ||||||||
-/comment | IAS | GWTC-3 | 4-OGC | IAS | GWTC-3 | 4-OGC | ||||
-GW190403_051519 | BBH_9 | |||||||||
-GW190408_181802 | BBH_2 | |||||||||
-GW190412_053044 | BBH_3 | |||||||||
-GW190413_052954 | BBH_6 | |||||||||
-GW190413_134308 | BBH_7 | |||||||||
-GW190421_213856 | BBH_7 | |||||||||
-GW190426_190642 | Veto (BBH_12) | 25.82 | 45.46 | |||||||
-GW190503_185404 | BBH_4 | |||||||||
-GW190512_180714 | BBH_1 | |||||||||
-GW190513_205428 | BBH_4 | |||||||||
-GW190514_065416 | BBH_7 | |||||||||
-GW190517_055101 | BBH_2 | |||||||||
-GW190519_153544 | BBH_7 | |||||||||
-GW190521_030229 | Veto (BBH_16) | 109.23 | 180.37 | |||||||
-GW190521_074359 | BBH_6 | |||||||||
-GW190527_092055 | BBH_6 | |||||||||
-GW190602_175927 | BBH_9 | |||||||||
-GW190701_203306 | Veto (BBH_7) | 34.87 | 55.9 | |||||||
-GW190706_222641 | BBH_8 | |||||||||
-GW190707_093326 | BBH_0 | |||||||||
-GW190719_215514 | BBH_8 | |||||||||
-GW190720_000836 | BBH_0 | |||||||||
-GW190725_174728 | Veto (BBH_3) | 27.35 | 55.99 | |||||||
-GW190727_060333 | BBH_6 | |||||||||
-GW190728_064510 | BBH_0 | |||||||||
-GW190731_140936 | BBH_6 | |||||||||
-GW190803_022701 | BBH_6 | |||||||||
-GW190805_211137 | BBH_6 | |||||||||
-GW190828_063405 | BBH_4 | |||||||||
-GW190828_065509 | Veto (BBH_1) | 50.14 | 56.48 | |||||||
-GW190909_114149111GW190909_114149 is a marginal candidate in GWTC-3 which has from our pipeline. | BBH_7 | |||||||||
-GW190915_235702 | BBH_4 | |||||||||
-GW190916_200658 | BBH_8 | |||||||||
-GW190917_114630 | Below | |||||||||
-GW190924_021846 | BBH_0 | |||||||||
-GW190926_050336 | BBH_6 | |||||||||
-GW190929_012149 | BBH_9 | |||||||||
-GW190930_133541 | Veto (BBH_1) | 36.02 | 45.98 | |||||||
-GW191103_012549 | BBH_0 | |||||||||
-GW191105_143521 | BBH_0 | |||||||||
-GW191109_010717 | Veto (BBH_8) | 82.74 | 164.37 |
-GWTC-3 events | Bank | IFAR (yr) | ||||||||
-/comment | IAS | GWTC-3 | 4-OGC | IAS | GWTC-3 | 4-OGC | ||||
-GW191113_071753 | Below | |||||||||
-GW191126_115259 | BBH_0 | |||||||||
-GW191127_050227 | BBH_8 | |||||||||
-GW191129_134029 | BBH_1 | |||||||||
-GW191204_110529 | BBH_2 | |||||||||
-GW191204_171526 | BBH_0 | |||||||||
-GW191215_223052 | BBH_2 | |||||||||
-GW191222_033537 | BBH_7 | |||||||||
-GW191230_180458 | BBH_8 | |||||||||
-GW200128_022011 | BBH_6 | |||||||||
-GW200129_065458 | Veto (BBH_4) | 203.73 | 342.75 | |||||||
-GW200202_154313 | BBH_1 | |||||||||
-GW200208_130117 | BBH_6 | |||||||||
-GW200208_222617 | BBH_8 | |||||||||
-GW200209_085452 | BBH_4 | |||||||||
-GW200210_092254 | Veto (BBH_3) | 25.53 | 48.5 | |||||||
-GW200216_220804 | BBH_8 | |||||||||
-GW200219_094415 | BBH_6 | |||||||||
-GW200220_061928 | Below | |||||||||
-GW200220_124850 | BBH_6 | |||||||||
-GW200224_222234 | BBH_6 | |||||||||
-GW200225_060421 | BBH_1 | |||||||||
-GW200306_093714 | BBH_5 | |||||||||
-GW200308_173609 | Below | |||||||||
-GW200311_115853 | BBH_4 | |||||||||
-GW200316_215756 | BBH_1 | |||||||||
-GW200322_091133 | BBH_6 | |||||||||
-4-OGC events | Bank | IFAR (yr) | ||||||||
-/comment | IAS | GWTC-3 | 4-OGC | IAS | GWTC-3 | 4-OGC | ||||
-GW191224_043228 | BBH_1 | 73.0 | ||||||||
-GW200106_134123 | BBH_6 | 55.0 | ||||||||
-GW200129_114245 | BBH_10 | 63.0 | ||||||||
-GW200210_005122 | Below | 70.0 | ||||||||
-GW200214_223306 | BBH_8 | 55.0 | ||||||||
-GW200305_084739 | Below | 59.0 | ||||||||
-GW200318_191337 | BBH_7 | 63.0 |
+| 1 | GW200104_184028 | 8,1 | | | | | ||||||||
+| 2 | GW190425_133124 | 14,0 | | | | | ||||||||
+| 3 | GW200319_022728 | 7,1 | | | | | ||||||||
+| 4 | GW190605_025957 | 14,0 | | | | | ||||||||
+| 5 | GW200114_020818 | 8,1 | | | | | ||||||||
+| 6 | GW190524_134109 | 9,0 | | | | | ||||||||
+| 7 | GW191213_204950 | 1,2 | | | | | ||||||||
+| 8 | GW190530_030659 | 3,2 | | | | | ||||||||
+| 9 | GW200210_035448 | 9,0 | | | | | ||||||||
+| 10 | GW190908_161547 | 7,1 | | | | | ||||||||
+| 11 | GW190710_150742 | 10,0 | | | | | ||||||||
 II.4 Astrophysical prior
@@
 where refer to detector-frame (redshifted) masses of the binary.
-We chose the location of the mass ratio cut for BBHs to restrict to the calibration regime of the IMRPhenomXAS approximant Pratten et al.
-(2020), while the cut on the individual spins is due to the relativistic Kerr limit.
-We picked a simple prior which broadly covers the parameter space (picking a broad prior can help with reweighting the results to a different prior in the future).
-Our prior on the mass ratio is a uniform one on for ; the effective spin parameter is flat within (with ); the total mass is distributed in a power-law fashion, .
+We chose the mass ratio cut for BBHs to restrict to the calibration regime of the IMRPhenomXAS approximant to numerical relativity simulations Pratten et al.
+(2020).
+We picked a simple prior which broadly covers the parameter space (picking a broad prior can help with reweighting the results to a different prior in the future Roulet et al.
+(2020)).
+We use a uniform prior on the mass ratio for ; the effective spin parameter is flat within (with ); the total mass is distributed in a power-law fashion, .
 Note that we use a prior on the detector-frame masses, and we do not assume a separate prior over redshift .
 We drew samples according to this prior (each sample corresponding to the physical parameter set ), and assigned a weight to each sample based on its observable volume.
-As a proxy for the observable distance, we simply calculated the SNR for the sample at a fixed distance by simulating its waveform using IMRPhenomXHM and using the same reference PSD which was used to make the template banks (weight SNR).
-Just for reference, for inspiral dominated waveforms, the observable volume roughly scales as Fishbach and Holz (2017).
-Using these samples, we then utilized the kernel density estimation method to construct probabilities for templates in our template banks (see appendix C of O22 for details on the density estimation method).
+We improve upon the astrophysical prior used in O22 by incorporating the fact that the observable changes across the parameter space.
+For example, for inspiral dominated waveforms, the observable volume roughly scales as Fishbach and Holz (2017).
+As a proxy for the observable distance, we simply calculated the SNR for the sample at a fixed distance by simulating its waveform using IMRPhenomXHM under a reference PSD (weightsample SNR3).
+Using these samples, we then utilized normalizing flows to predict the effective astrophysical prior density at the location of our templates.
+We also use normalizing flows to generate physical and samples corresponding to each template (which are required by our marginalization integral).
+We leave further discussion of our use of normalizing flows to an upcoming paper Wadekar et al.
+(b).
 Fig.
 7 in the Appendix shows the cumulative probabilities for our banks, each of which is obtained by adding the weights of all the samples falling within the corresponding bank.
-II.4.1 Comparison with astrophysical prior used in our -only banks
-It is worth briefly comparing the astrophysical distributions we assumed here to those we assumed in our previous -only O3 searches.
-We split our banks roughly according to the detector-frame chirp mass (the bank edges were uniform in log ) and further split each bank into subbanks based on the amplitude profiles of normalized waveforms Roulet et al.
-(2019).
-We assumed that the total prior astrophysical probability within each of our banks was identical, and that the probability distribution within each bank was flat in the detector-frame component masses and also in the effective spin (without taking the observable distance into account; we take this effect into account in the current search).
-Note that we also slightly increased our search space in this paper: previously, our banks were limited to and (see figure 2 of O22), while we now use .
 III Results
 III.1 New candidate events
-Using our new search pipeline on the Hanford–Livingston data from the third observing run (O3), we find 14 candidate events with which have not been previously reported in the literature.
+Using our new search pipeline on the Hanford–Livingston data from the third observing run (O3), we find 11 candidate events with which have not been previously reported in the literature.
 We list the properties of these new candidates in Table 1.
@@
 Note that the high-mass banks in our case have much fewer templates than the low-mass banks (see Table I of Wadekar et al.
-(2023a)), which leads to considerable heterogeneity in the number of background triggers in the banks.
+(2023)), which leads to considerable heterogeneity in the number of background triggers in the banks.
 Hence, the values needed for events to be above threshold can be different for different banks.
@@
 We estimate the values in a manner that is similar to that outlined in Appendix B of O22, from the distribution of foreground and background triggers according to
-where in the above formula corresponds to our final ranking statistic value for the candidate events (see Ref.
-Wadekar et al.
-(2023b) for the full expression), and corresponds to number density values from a 1D histogram of the ranking statistic (which we separately calculate for the signal and background cases).
+where in the above formula corresponds to our final ranking statistic value for the candidate events (see Eq.
+16 of Ref.
+W24 for the full expression), and corresponds to number density values from a 1D histogram of the ranking statistic (which we separately calculate for the signal and background cases).
 Note that as the templates in our banks have astrophysical probabilities associated with them (see Section II.4), and we use these values in calculating .
@@
 We expect this dependence to be more significant for the ‘overall’ IFAR in comparison to the ‘per bank’ IFAR values quoted in Table 1.
-We therefore separately provide the ‘per bank’ values in our table, so one can reweight them in a rough way for different astrophysical distributions than the ones assumed here.
+We therefore separately provide the ‘per bank’ values in our table, so one could reweight them in a rough way for different astrophysical distributions than the ones assumed here.
+| Event | Bank | PE (90% credibility) | (evidence) | IFAR (yr) | |||||||||
+| HM | HM+Prec. | per bank | overall | ||||||||||
+| GW190711_030756 | 6,2 | | | | | | | | |||||
+| GW200109_195634 | 9,0 | | | | | | | | |||||
+| GW190818_232544 | 8,1 | | | | | | | | |||||
+| GW190707_083226 | 7,1 | | | | | | | | |||||
+| GW191117_023843 | 10,0 | | | | | | | | |||||
+| GWTC-3 events | Bank | GWTC-3 | IFAR (yr) | |||||||
+| /comment | IAS | GWTC-3 | 4-OGC | IAS | GWTC-3 | 4-OGC | ||||
+| GW190403_051519 | BBH_6 | |||||||||
+| GW190408_181802 | BBH_2 | |||||||||
+| GW190412_053044 | BBH_1 | |||||||||
+| GW190413_052954 | BBH_5 | |||||||||
+| GW190413_134308 | BBH_7 | |||||||||
+| GW190421_213856 | BBH_5 | |||||||||
+| GW190426_190642 | Veto (BBH_12) | 20.88 | 58.34 | |||||||
+| GW190503_185404 | BBH_5 | |||||||||
+| GW190512_180714 | BBH_1 | |||||||||
+| GW190513_205428 | BBH_3 | |||||||||
+| GW190514_065416 | BBH_6 | |||||||||
+| GW190517_055101 | BBH_3 | |||||||||
+| GW190519_153544 | BBH_7 | |||||||||
+| GW190521_030229 | BBH_11 | |||||||||
+| GW190521_074359 | BBH_5 | |||||||||
+| GW190527_092055 | BBH_5 | |||||||||
+| GW190602_175927 | BBH_8 | |||||||||
+| GW190701_203306 | Veto (BBH_7) | 33.32 | 53.2 | |||||||
+| GW190706_222641 | BBH_8 | |||||||||
+| GW190707_093326 | BBH_0 | |||||||||
+| GW190719_215514 | BBH_5 | |||||||||
+| GW190720_000836 | BBH_0 | |||||||||
+| GW190725_174728 | BBH_0 | |||||||||
+| GW190727_060333 | BBH_5 | |||||||||
+| GW190728_064510 | BBH_0 | |||||||||
+| GW190731_140936 | BBH_6 | |||||||||
+| GW190803_022701 | BBH_5 | |||||||||
+| GW190805_211137 | BBH_6 | |||||||||
+| GW190828_063405 | BBH_3 | |||||||||
+| GW190828_065509 | BBH_1 | |||||||||
+| GW190909_114149222GW190909_114149 is a marginal candidate in GWTC-3 which has from our pipeline. | BBH_5 | |||||||||
+| GW190915_235702 | BBH_3 | |||||||||
+| GW190916_200658 | BBH_6 | |||||||||
+| GW190917_114630 | BBH_0 | |||||||||
+| GW190924_021846 | BBH_0 | |||||||||
+| GW190926_050336 | BBH_6 | |||||||||
+| GW190929_012149 | BBH_8 | |||||||||
+| GW190930_133541 | BBH_0 | |||||||||
+| GW191103_012549 | BBH_0 | |||||||||
+| GW191105_143521 | BBH_0 | |||||||||
+| GW191109_010717 | BBH_7 | |||||||||
+| GWTC-3 events | Bank | GWTC-3 | IFAR (yr) | |||||||
+| /comment | IAS | GWTC-3 | 4-OGC | IAS | GWTC-3 | 4-OGC | ||||
+| GW191113_071753 | BBH_1 | |||||||||
+| GW191126_115259 | BBH_0 | |||||||||
+| GW191127_050227 | BBH_7 | |||||||||
+| GW191129_134029 | BBH_0 | |||||||||
+| GW191204_110529 | BBH_2 | |||||||||
+| GW191204_171526 | BBH_0 | |||||||||
+| GW191215_223052 | BBH_2 | |||||||||
+| GW191222_033537 | BBH_6 | |||||||||
+| GW191230_180458 | BBH_7 | |||||||||
+| GW200128_022011 | BBH_6 | |||||||||
+| GW200129_065458 | BBH_3 | |||||||||
+| GW200202_154313 | BBH_1 | |||||||||
+| GW200208_130117 | BBH_5 | |||||||||
+| GW200208_222617 | BBH_6 | |||||||||
+| GW200209_085452 | BBH_5 | |||||||||
+| GW200210_092254 | Veto (BBH_1) | 29.3 | 45.4 | |||||||
+| GW200216_220804 | BBH_7 | |||||||||
+| GW200219_094415 | BBH_5 | |||||||||
+| GW200220_061928 | BBH_11 | |||||||||
+| GW200220_124850 | BBH_6 | |||||||||
+| GW200224_222234 | BBH_5 | |||||||||
+| GW200225_060421 | BBH_1 | |||||||||
+| GW200306_093714 | BBH_2 | |||||||||
+| GW200308_173609 | Below | |||||||||
+| GW200311_115853 | BBH_3 | |||||||||
+| GW200316_215756 | BBH_1 | |||||||||
+| GW200322_091133 | BBH_1 | |||||||||
+| 4-OGC events | Bank | 4-OGC | IFAR (yr) | |||||||
+| /comment | IAS | GWTC-3 | 4-OGC | IAS | GWTC-3 | 4-OGC | ||||
+| GW191224_043228 | BBH_0 | 73.0 | ||||||||
+| GW200106_134123 | BBH_6 | 55.0 | ||||||||
+| GW200129_114245 | BBH_7 | 63.0 | ||||||||
+| GW200210_005122 | BBH_1 | 32.4 | 42.1 | 70.0 | ||||||
+| GW200214_223306 | BBH_7 | 55.0 | ||||||||
+| GW200305_084739 | Below | 59.0 | ||||||||
+| GW200318_191337 | BBH_7 | 63.0 | ||||||||
 III.1.1 Parameter estimation runs
-We perform parameter estimation (PE) runs using the package222https://github.com/jroulet/cogwheel Roulet et al.
+We perform parameter estimation (PE) runs using the package333https://github.com/jroulet/cogwheel Roulet et al.
 (2022) with the sampler Lange (2023).
@@
 Zackay et al.
+(2018); Roulet et al.
 (2024) as implemented in cogwheel.
@@
 This helps us calculate the relative evidence (Bayes factor) due to including the effects of HM and HM + precession separately for the new candidate events; we show the logarithm of the Bayes factors in Table 1.
-Four of the new candidate events (the ones numbered ) have Bayes factors larger than unity for HM, and the detection of these candidates likely benefited the most from our search containing HM.
-Most of the remaining events have Bayes factors for HM that are lower than unity, and we judge that the detection of these additional candidates was a result of just an overall improvement of our search’s sensitivity in the high mass/redshift region due to e.g., improved template coverage and background estimation.
+Two of the new candidate events have Bayes factors larger than unity for HM, and the detection of these candidates likely benefited the most from our search containing HM.
 We show a few select 2D posteriors of our new candidate events in colored contours in Figs.
-1, 2 and 7.
+1, 2 and 6.
 We also overplot the contours from events in GWTC-3, 4-OGC and the IAS -only O3 catalogs in light gray.
 Our search only covers events with Hanford–Livingston coincidence triggers, but we include all binary BH events that are part of the LVK catalogs through O1–O3 in these figures.
-Note that the posteriors are bi-modal for some events, which show up as two contours in the plot.
 We add dashed lines to mark the approximate limits of sensitivity in the top panel of Fig.
 1.
-The dashed line in the top-right of the panel marks the region where the binary starts moving out of band (we take this to happen when the detector frame mass equals , which is the limit of our template bank).
+The dashed line in the top-right of the panel marks the region where the binary starts moving out of band (we roughly take this to happen when the detector frame mass equals , which is the limit of our template bank).
 The dashed line in the top-left marks the contour of constant detector-frame luminosity for an inspiral-dominated equal-mass binary; it roughly delineates the parameter limits past which binaries are too faint to be observable.
@@
 Figs.
-10 and 11 in the Appendix show full corner plots from different PE runs.
-Some of the events in Fig.
-3 (e.g., the ones numbered ) have posteriors that show significant support for large, positive effective spin under the uniform in prior.
-The isotropic spin prior recovers a lower likelihood for these events, since the prior pulls the sampler away from the high solution.
-GW191228_195619 has the largest evidence for HM () and also for the combination of HM+precession ().
-We show a corner plot corresponding to this event in Fig.
-4 for our three PE runs, and we indeed see significant support for both HM and precession.
-Note that for the (green) and + HM (orange) cases, the distribution of the primary component’s in-plane spin component is that of its prior conditioned on the other parameters.
-For these cases, we set the in-plane spin components to zero when evaluating likelihoods.
-Within 90% CL, we find to be .
-Note also that the total SNR (110) recovered by our non-precessing Hanford+Livingston search is significantly smaller than obtained from a PE run which includes both the effects of orbital precession and data from the Virgo detector.
-Therefore, we conjecture that the significance of this event could be improved in future searches that include the effects of precession and/or Virgo data.
+9 and 10 in the Appendix show full corner plots from different PE runs.
+We additionally highlight the posteriors for two most interesting new candidate events (GW200114_020818, GW190425_133124) in Fig.
+5.
+GW200114_020818 has the largest evidence for HM ().
+We show three PE runs with and without HM and precessing effects.
+Note that for the non-precessing (green and orange) cases, the distribution of the in-plane spin component is same as that of the prior conditioned on the other parameters (we set contribution of in-plane spin components to be zero inside the likelihood for these cases).
+Note also that the total SNR2 (180) recovered by our non-precessing Hanford+Livingston search is significantly smaller than obtained from a PE run which includes both the effects of orbital precession and data from the Virgo detector.
+Similarly, SNR77 and for GW190425_133124.
+Therefore, we predict that the significance of these event could be improved in future searches that include the effects of precession and/or Virgo data.
 We also compare our results against PE runs using the IMRPhenomXPHM approximant Pratten et al.
-(2021) (under the two different MSA Chatziioannou et al.
-(2017) and NNLO Hannam et al.
-(2014) prescriptions for evolving the spins) in Fig.
+(2021) (using the default MSA Chatziioannou et al.
+(2017) prescription for evolving the spins) in Fig.
 8 in the Appendix; we find good agreement in most of the cases.
@@
 The full comparison catalog including lower candidates is given in Table 6.
-Overall, we recover 6 events with .
-We caution against directly comparing the sensitivities of the -only and + HM pipelines from Table 6 as there are multiple differences: () including HM in the search, () using a different astrophysical prior (see Section II.4) () additional glitch cleaning at high masses (see Section II.3).
+Overall, we recover 5 events with .
+We caution against directly comparing the sensitivities of the -only and + HM pipelines from Table 6 as there are multiple differences: () including HM in the search, () using a different astrophysical prior (see Section II.4) () additional glitch mitigation at high masses (see Section II.3).
 It is worth mentioning though that we recover the events which were reported in previous IAS catalogs with , except GW190704_104834 (this event was in BBH-0 in O22 and we checked that the difference in the astrophysical priors between the pipelines was large for that bank).
-We collect a larger number of triggers in the HM search as compared to a -only search, so we raised the collection threshold to mitigate the increased memory requirements.
+We collect a larger number of triggers in the HM search as compared to a -only search, so we raised the collection threshold to mitigate the increased memory requirements (see Section II.2.1).
 This caused us to miss low-SNR events that were reported by -only pipelines with lower collection thresholds in the previous literature.
-We further discuss this point in Section III.3.2 below.
 One notable event in Table 2 is GW190711_030756.
-The IFAR of this event in our HM search is 38.7 yr as opposed to yr in the -only search O22 (after accounting for the trials factor in O22 due to the number of banks searched).
+The IFAR of this event in our HM search is 11 yr as opposed to yr in the -only search O22 (after accounting for the trials factor in O22 due to the number of banks searched).
 We show a corner plot for this event in Fig.
-5.
-Indeed, GW190711_030756 also shows a significant evidence for higher modes and our search recovers an additional 10 points in SNR compared to the -only search.
+4.
+Indeed, GW190711_030756 also shows a significant evidence for higher modes and our search recovers an additional 10 points in SNR2 compared to the -only search.
 See Fig.
@@
 III.3 Comparison to the LVK and OGC -only catalogs
-In Table 3, we compare our pipeline’s results for the O3 Hanford–Livingston coincident events published in the GWTC-3 catalog333https://gwosc.org/GWTC-3/ Abbott et al.
+In Table LABEL:tab:lvc_events, we compare our pipeline’s results for the O3 Hanford–Livingston coincident events published in the GWTC-3 catalog444https://gwosc.org/GWTC-3/ Abbott et al.
 (2021).
@@
 (2016).
-We also include the comparison with the 4-OGC catalog taken from the summary data in the repository444https://github.com/gwastro/4-ogc/blob/master/search/4OGC_top.txt listed in Nitz et al.
+We also include the comparison with the 4-OGC catalog taken from the summary data in the repository555https://github.com/gwastro/4-ogc/blob/master/search/4OGC_top.txt listed in Nitz et al.
 (2021).
 Note that the quoted GWTC-3 squared SNR, , is from LVK’s PE runs for the events.
-The search SNR is generally lower, and varies between the different pipelines that the LVK uses.
+The search SNR2 is generally lower, and varies between the different pipelines that the LVK uses.
 We broadly recover the confident LVK detections, except some vetoed/missed events which are discussed in more detail in the subsections below.
-Overall, 9 events GWTC-3 events in Table 1 were vetoed in our pipeline.
-4 GWTC-3 events fell below our trigger collection SNR threshold (GW190917_114630, GW191113_071753, GW200220_061928, GW200308_173609), note that these were confidently detected by only one of the LVK pipelines.
-Three GWTC-3 events were detected with in our pipeline (GW200208_222617, GW200306_093714, GW200322_091133), note again that these were confidently detected by only one of the LVK pipelines.
-At the end of Table 3, we include comparison with the list of 7 events in 4-OGC which are not present in GWTC-3 catalog (from which we recover 4 with and 2 were below our collection SNR threshold).
+Overall, three GWTC-3 events in Table 1 were vetoed in our pipeline.
+GW190701_203306 was vetoed due to excess power present after subtracting the best-fit waveform in Livingston (note that the data around this event was treated for scattered light glitches Abbott et al.
+(2021)).
+GW200308_173609 has GWTC-3 and it fell below our SNR collection threshold.
+Six GWTC-3 events were detected with in our pipeline (GW190403_051519, GW190917_114630, GW191113_071753, GW200202_154313, GW200208_222617, GW200322_091133).
+Note that all the GWTC-3 events we missed (except GW200202_154313) were confidently detected by only one of the LVK pipelines.
+We find GW200202_154313 gets penalized in the multi-detector coherence module of our pipeline possibly due to having a large SNR difference in the two detectors.
+At the end of Table LABEL:tab:lvc_events, we include comparison with the list of 7 events in 4-OGC which are not present in GWTC-3 catalog (from which we recover 5 with ).
 GW190909_114149 is a marginal candidate in GWTC-3 which has from our pipeline.
-We still include it in Table 3 as it was first a significant candidate in GWTC-2 Abbott et al.
-(2021), but later updated to be in the marginal list The LIGO Scientific Collaboration et al.
+We still include it in Table LABEL:tab:lvc_events as it was first a significant candidate in GWTC-2 Abbott et al.
+(2021), but later updated to be in their marginal list The LIGO Scientific Collaboration et al.
 (2021).
-We also checked the overlap of our new candidate event list with the sub-threshold candidate list (i.e., which have but FAR day) released by the LVK collaboration: GWTC-2.1 and GWTC-3 The LIGO Scientific Collaboration et al.
+We also checked the overlap of our new candidate event list with the sub-threshold candidate list (i.e., which have but FAR day-1) released by the LVK collaboration: GWTC-2.1 and GWTC-3 The LIGO Scientific Collaboration et al.
 (2021); Abbott et al.
@@
 (2023).
-III.3.1 Vetoed events
-Our pipeline includes a series of signal consistency checks in order to veto triggers that are likely to arise from noise transients (glitches) to improve our search sensitivity (see section I of Venumadhav et al.
-(2019)).
-The vetoes are designed in a way that they are falsely triggered by real signals with only a low probability (a few percent) if their waveforms are accurately modeled by templates in the bank, and the noise is Gaussian.
-If a loud signal displays effects which are not present in our templates, like precession or eccentricity, the chances of the veto mis-identifying it as a glitch can be much higher.
-GW190521_030229, GW191109_010717 and GW200129_065458 were reported to be fairly high SNR with imprints of precession Abbott et al.
-(2020, 2021), which could be the reason for them being vetoed.
-For example, the presence of higher modes in GW190521 is disfavored with a Bayes factor of , while a precessing signal is favored with a factor of Abbott et al.
-(2020).
-GW190701_203306 was vetoed due to excess power present after subtracting the best-fit waveform (note that the data around this event was treated for scattered light glitches Abbott et al.
-(2021)).
-Out of the veto tests that we apply to events, we noticed GW190828_065509 and GW190930_133541 each failed just one test (namely, the split test with chunk sets, and the test with 20 chunks, respectively).
-We noticed a small issue in the pipeline where we artificially veto triggers very close ( s) to the edge of the timeseries files; this led to the event GW190725_174728 being vetoed.
-We will fix this issue in future searches.
-The event GW190426_190642 was detected confidently by only one of the LVK search pipelines and got vetoed in our case.
-We expect to recover some of loud vetoed events if we also include precession in our template banks.
-Alternatively, one could formulate a procedure to override vetos or changing their thresholds based on injections with precessing waveforms.
-We leave exploring these directions to future work.
-III.3.2 Other missed/downweighted events
 As we discussed earlier in Section II.2, we set a particular SNR threshold () to collect triggers from individual detectors.
-We do this in order to regulate the number of single-detector triggers that are passed to the more costly coincidence step of the pipeline.
-If we set the threshold based on the quadrature sum of the SNRs of the (orthogonalized) harmonics, we collect more triggers in a HM search as compared to a -only search with the same threshold.
-Hence, we raised the collection threshold (see Section II.2.1 for a detailed discussion).
+We do this in order to regulate the number of single-detector triggers that are passed to the more costly coincidence step of the pipeline (see section II.2.1 for a detailed discussion).
 Due to this, we missed low-SNR events that were reported by -only pipelines in the previous literature.
-We label the events which fell below this threshold as “Below ” in the table, where the threshold depends on the number of templates in the bank and is therefore different for different banks.
-On further investigation, using a smaller value of is feasible with our current search setup, and thus we plan to lower the threshold in future searches.
-We do not include the following events in Table 3 because they were detected in Livingston–Virgo or Hanford–Virgo coincidence, or single detector search, all of which we have yet to run: GW190620_030421, GW190630_185205, GW190708_232457, GW190814, GW190910_112807, and GW190925_232845.
+We label the events which fell below this threshold as “Below ” in the table.
+We do not include the following events in Table LABEL:tab:lvc_events because they were detected in Livingston–Virgo or Hanford–Virgo coincidence, or single detector search, all of which we did not run: GW190620_030421, GW190630_185205, GW190708_232457, GW190814, GW190910_112807, and GW190925_232845.
 We also only searched for BBH mergers in this paper, and did not recover the NSBH candidates GW191219_163120 and GW200115_042309 (in addition, the posterior probability for GW200210_092254 peaks in the NSBH region Abbott et al.
-(2021) and thus it was vetoed in our search).
+(2021), which could lead to it being vetoed in our search).
 We hope to perform coincident searches which would also include Virgo (and KAGRA) in the future.
 We also intend to perform a separate search for NSBH mergers including our HM pipeline in a future paper.
-GW191222_033537 was detected with in our pipeline but has comparatively a much higher false alarm rate.
-We checked that it was downweighted due to being in the proximity of a glitch (which led to the event acquiring unphysical amounts of SNR in HM).
 IV Interesting astrophysical properties of the new candidate events
-In this section, we discuss interesting astrophysical properties of the 14 new candidate events, whose parameters we have reported in Table 1 and in Figs. 1, 2 and 3.
+In this section, we discuss interesting astrophysical properties of the 11 new candidate events, whose parameters we have reported in Table 1 and in Figs. 1, 2 and 3.
 IV.1 Masses in the IMBH range and pair instability mass gap
-Black holes roughly within the mass range – are designated as intermediate-mass black holes (IMBHs).
+BHs roughly within the mass range – are designated as intermediate-mass black holes (IMBHs).
+There are a number of channels which can produce BHs in this mass regime, see e.g., Coleman Miller and Hamilton (2002); Franciolini et al.
+(2024).
 Several IMBH candidates have been suggested by electromagnetic observations, but these lack conclusive confirmation Greene et al.
@@
 (2018).
-In our search, at % credibility, 4 candidate events have total masses in the IMBH range (GW190605_025957, GW190604_103812, GW191228_195619 and GW190530_133833).
-Among them, GW191228_195619 has the highest total mass (and also has the highest primary mass ), which makes it, by a large margin, the most massive candidate with detected in gravitational wave data so far.
+In our search, at % credibility, 4 candidate events have primary black hole masses in the IMBH range (# 2, 4, 5, 9 in Table 1).
 Another interesting regime in BH masses is the pair instability mass gap.
@@
 These processes are expected to produce a gap roughly between – in the distribution of BH mass.
+Many of our new candidate events have support for masses in the gap.
 Note however that the exact boundaries of such a gap are not well determined, as they depend on multiple uncertain factors such as the rate of carbon to oxygen burning in the core , angular momentum transport, metallicity, wind loss, etc.
 Marchant and Moriya (2020); Farmer et al.
+(2019); van Son et al.
 (2020); Mehta et al.
 (2022); Hendriks et al.
+(2023); Golomb et al.
+(2023); Franciolini et al.
 (2024).
@@
 (2021); Fragione et al.
-(2021), or from the gaseous disk channel Tagawa et al.
+(2021), stellar collisions Kremer et al.
+(2020), or from the gaseous disk channel Yang et al.
+(2019); McKernan et al.
+(2020); Tagawa et al.
+(2020); Ishibashi, W.
+and Gröbner, M.
+(2020); Secunda et al.
+(2020); Tagawa et al.
+(2021); Fabj et al.
+(2020); Tagawa et al.
 (2020); Samsing et al.
 (2022).
-In our search, at 90% CL, 9 candidates have median mass of the primary BH lying between – (# 1, 4, 5, 6, 8, 10, 11, 13, 14 in Table 1).
-One of our candidates: GW191228_195619 has its secondary BH mass overlapping with the mass-gap range.
 Our new candidate events could be helpful in estimating the abundance of BH (or rate of BH mergers) in/beyond the pair-instability mass regime.
 This can also shed light on BH progenitor channels.
+Other potential explanation for high-mass merger detections could be that they are lower-mass mergers but are lensed by a foreground galaxy Dai et al.
+(2017).
 IV.2 High redshifts
-In our search, we find that 9 of our new candidates have redshift at % CL (# 1, 2, 3, 6, 8, 10, 11, 13, 14 in Table 1).
+We find two new candidates with redshift at % CL (# 4, 6 in Table 1).
+5 of our new candidates have median redshift .
 Finding high-redshift BHs can be important to answer questions such as: How does the merger rate of BHs evolve with redshift Fishbach et al.
@@
 Does the merger rate change in a different manner compared to the the star formation rate (which peaks at Madau and Dickinson (2014))?
-Furthermore, different BH formation channels can have different predictions for the merger rate at high redshift Mapelli et al.
+Furthermore, different BH formation channels have different predictions for the merger rate at high redshift Mapelli et al.
 (2022); van Son et al.
+(2022); Boesky et al.
+(2024a, b).
+For example, the chemically homogeneous evolution channel or the stable mass transfer channel produces BBHs of relatively higher masses and lower redshifts than the common envelope evolution channel Mandel and de Mink (2016); van Son et al.
 (2022).
-For example, the chemically homogeneous evolution channel produces BBHs of relatively higher masses and lower redshifts than the common envelope evolution channel Mandel and de Mink (2016).
 In the dynamical channel, massive BHs can have relatively longer delay times compared to star formation Mapelli et al.
@@
 We aim to test some of these models in a future population inference study.
-IV.3 Preference for positive effective spins
-Measuring spins of the black holes is one of the best ways to distinguish between isolated and dynamical formation channels.
+IV.3 Preference for non-zero spins
+Measuring spins of the black holes in a binary is one of the best ways to distinguish between isolated and dynamical formation channels.
 In the isolated channel, accretion can lead to spins of BHs which are aligned in the direction of the binary angular momentum (leading to positive values of the effective spin parameter ).
 Dynamical channels, on the other hand, predict the astrophysical distribution to be symmetric around 0.
-This is because the the spins BH captured into a binary are not expected to be preferentially aligned along any particular direction.
-From the PE runs under our fiducial prior, 6 candidates have effective spin at credibility (# 2, 4, 6, 8, 9, 11 in Table 1).
-Note however that posteriors are significantly dependent on the prior as seen in the comparison in Fig.
-3.
-Though, even under the isotropic spin prior, some of the new candidate events still have mild preference for .
-Note that the preference for in our detection catalog does not immediately mean that astrophysics prefers this scenario, as this could be caused by selection effects.
-One of the selection effects is that positive binaries are louder than their negative counterparts; due to the orbital hangup effect, the positive binaries merge later and have power at higher frequencies Ng et al.
-(2018); Roulet and Zaldarriaga (2019).
-This effect is much more prominent for high mass binaries (where a significant portion of signal is masked by the low frequency end of the PSD) as we show in Fig.
-6.
-There is another effect which further downweights high-mass negative signals.
-Their waveforms have a short duration in band and therefore can be more effectively mimicked by glitches.
-This degrades the sensitivity to such systems in our search Wadekar et al.
-(2023b) when we perform the background dependent reweighting in our ranking statistic (see Section II.3.3 and Ref.
-Wadekar et al.
-(2023b)).
-To robustly determine if the true astrophysical distribution also favors at high masses, we plan to a do a population study (where we will use injections with both positive and negative values to determine our pipeline selection function).
+This is because the spins BH captured into a binary are not expected to be preferentially aligned along any particular direction.
+From the PE runs under our fiducial prior (flat in ), two new candidates have positive effective spin at CL (GW191213_204950, GW190530_030659).
+On the other hand, GW200114_020818 has at CL.
+Even under the isotropic spin prior, some of the new candidate events still have mild preference for being non-zero.
 IV.4 Asymmetric mass ratio systems
-Five of our new candidates have median mass ratio (# 4, 6, 7, 10, 11 in Table 1).
+5 candidates have median mass ratio .
+The posteriors are broad in most cases, but there are two candidates with a strong preference for asymmetric mass-ratios (GW200114_020818 and GW200210_035448).
 Finding sources with asymmetric mass ratios can shed light on the formation channels Abbott et al.
@@
 This is because higher mass BHs in a globular cluster sink to their cores due to dynamical friction and vice versa for lighter BHs.
-The caveat is that if clusters can retain merger products, they can produce “second-generation” mergers whose mass ratios can deviate significantly from unity Rodriguez et al.
+Although there is a caveat here that if clusters can retain merger products, they can produce “second-generation” mergers whose mass ratios can deviate significantly from unity Rodriguez et al.
 (2019).
 In the isolated binary channel, in certain cases, an initially asymmetric mass ratio system can evolve toward a more symmetric configuration due to mass accretion episodes Ivanova et al.
-(2013); Postnov and Yungelson (2014); note that the reverse scenario has been proposed as well (see e.g., Ref.
-Broekgaarden et al.
+(2013); Postnov and Yungelson (2014); note that the reverse scenario has been proposed as well (see e.g., Gerosa et al.
+(2013); Broekgaarden et al.
 (2022)).
@@
 V Discussion
-V.1 Dependence on the choice of astrophysical prior
+V.1 Injection campaigns to quantify gain in sensitivity due to including HM
+We quantify the improvement in the sensitive volume-time () of the search pipeline in our companion paper: Mehta et al.
+(2025).
+We use the injection catalog released by the LVK collaboration on Zenodo LIGO Scientific Collaboration (2021).
+We also compare the results of our IAS-HM pipeline with its (2,2)-only limit.
+We find significant improvement in for systems with asymmetric mass ratios and larger total masses, with gains in sensitivity even exceeding 100% at certain high masses (see Fig.
+2 of Mehta et al.
+(2025)).
+We also do not find a loss of sensitivity in the IAS-HM pipeline for near equal mass-ratio mergers.
+We also find that the of the IAS pipeline is either larger than or comparable to that of the individual LVK pipelines (see Fig.
+1 of Mehta et al.
+(2025)).
+It is quite encouraging to see that the parameter space where we get a large improvement () over the LVK and the (2,2)-only IAS pipelines is similar to the parameter space where we detect most of the new candidate events, thus adding more credibility to our newly reported detections.
+We refer the reader to Mehta et al.
+(2025) for further details on the injection campaigns.
+Previous studies have noted that the larger variety in waveform morphologies (brought by the inclusion of higher modes) can lead to a loss of sensitivity in the search, as the number of background triggers is increased, see e.g., Capano et al.
+(2014); Chandra et al.
+(2022).
+This can be understood from the fact that, with HMs, the shape of the waveform also depends on the inclination and orbital phase, and these additional degrees of freedom increase the trials factor when finding the best-matching template.
+As discussed in section II.2 and W24, we do not naively add additional degrees of freedom to the templates throughout the parameter space.
+Instead, our new detection statistic (see Section II.2 and W24) includes HM while simultaneously marginalizing the likelihood over the extrinsic parameters and mass ratio.
+This marginalization naturally downweights fine-tuned solutions, approaching the optimal statistic and preventing an undue increase of the trials factor (e.g., the edge-on solutions get naturally downweighted when we marginalize over distance, as their observable volume is smaller).
+The downside of using a prior distribution on is that it is partly dependent on our choice of astrophysical prior, which is currently not precisely known.
+Especially, the mass ratio distribution is not very well constrained and affects our prior for .
+Using a different distribution for than the true astrophysical distribution can indeed lead to some loss of sensitivity.
+We chose a broad prior (flat in ) in our search, with the motivation that such a prior can be easily reweighted later in a hierarchical population analysis.
+Even though including HM can cause some loss in sensitivity for near-threshold face-on equal mass systems, note that detecting even a few events with asymmetric-mass ratios can inform our population estimates much more than detecting a slightly larger number of equal-mass systems.
+It is worth drawing an analogy here to the case of spin distribution of templates used in -only searches.
+Neglecting templates with negative indeed improves the search sensitivity to near-threshold systems, but finding even a few systems can greatly help us distinguish between progenitor channels of BHs.
+We thus argue the goal of a search should not solely be to maximize the total number of detections, but to holistically cover most of the interesting regions of the binary parameter space.
+A similar argument also holds for when including spin-precession or eccentricity effects in the search templates.
+Detecting even a few events with significantly misaligned spins can shed more light on the formation channels of binaries as compared to detecting a larger number of near-threshold aligned/zero spin mergers.
+Some of the challenges for including these effects are similar to the ones for HM: significant increase in the template bank size and loss of sensitivity due to collecting additional background McIsaac et al.
+(2023); Schmidt et al.
+(2024a, b); Dhurkunde and Nitz (2022, 2023); Nitz et al.
+(2020).
+We will develop methods similar to the ones implemented in our search to help tackle those searches in future work Zhou et al.
+; Islam et al.
+.
+V.2 Dependence on the choice of astrophysical prior
 Our estimate of and our overall IFAR depend on the astrophysical prior used in the search, and that dependence is strongest in the case of near-threshold detections like some of the new candidates in our catalog list in Table 1.
-To facilitate further analysis and comparisons between different astrophysical priors, we make our posterior samples and other relevant information public at https://github.com/JayWadekar/GW_higher_harmonics_search.
+To facilitate further analysis and comparisons between different astrophysical priors, we make our posterior samples, the selection function of our pipeline and other relevant information public at https://github.com/JayWadekar/GW_higher_harmonics_search.
 These can be used to reweight our results with other choices of priors (see e.g., Payne et al.
@@
 Another difference in our priors is for the case of total mass and mass ratios.
-It is also worth discussing how our other priors compare with the results of Abbott et al.
-(2023).
 We used a search prior uniform in , whereas, Ref.
@@
 A full population analysis not only requires the PE samples but also the of each event Roulet et al.
+(2020); Gaebel et al.
 (2019).
-For this reason, we also provide a file, IAS_HM_O3_triggers.hdf, in the repository that contains the full list of triggers (including the sub-threshold ones).
-Just for reference, we report subthreshold events () in Table 7 in the Appendix.
-V.2 What is the gain/loss in sensitivity due to including HM in the search?
-Unfortunately, we cannot currently answer the above question by comparing the results from our search to those from the previous IAS -only searches Olsen et al.
-(2022); Mehta et al.
+For this reason, we also provide a file, IAS_HM_O3_triggers.json, in the repository that contains the full list of triggers (including the sub-threshold ones).
+We also include the subthreshold events () in Table 7 in the Appendix for reference.
+V.3 Population study
+To probe interesting population inference questions in section IV, one needs to perform a hierarchical bayesian analysis and account for selection function of our search.
+We leave this analysis to an upcoming work Mehta et al.
+.
+The new candidate events reported here have moderate IFARs (the best event has yr), thus a methodology to combine confident and marginal events in population inference will be required Roulet et al.
 (2020).
-This is because our search not only includes the effect of HM, but also includes the improvements (for all modes) due to downweighting noise transients (see Section II.3 and Wadekar et al.
-(2023b)).
-Furthermore, compared to earlier, we use a different astrophysical prior (see section II.4) and use a different waveform approximant for modeling the waveforms (IMRPhenomXAS Pratten et al.
-(2020) instead of IMRPhenomD Khan et al.
-(2016)).
-For us to isolate the effect of HM, we will need to rerun our search (ideally on injections) with all the settings to be the same except having only modes in the template banks.
-We leave this study to a future work.
-Previous studies have noted that the larger variety in waveform morphologies (brought by the inclusion of higher modes) can lead to a loss of sensitivity in the search, as the number of background triggers is increased, see e.g., Capano et al.
-(2014); Chandra et al.
-(2022).
-This can be understood from the fact that, with higher modes, the shape of the waveform also depends on the inclination and orbital phase, and these additional degrees of freedom increase the trials factor when finding the best-matching template.
-In our search, we do not blankly add additional degrees of freedom to the templates throughout the parameter space.
-Instead, our new detection statistic (see Section II.2 and Wadekar et al.
-(2023b)) includes HM while simultaneously marginalizing the likelihood over the extrinsic parameters (which are the same as in the -only search).
-This marginalization naturally downweights fine-tuned solutions, approaching the optimal statistic and preventing an undue increase of the trials factor (e.g., the edge-on solutions get naturally downweighted when we marginalize over distance, as their observable volume is smaller).
-In our ranking statistic for the triggers, we also marginalize over the ratios of the SNR amplitudes of HM: and .
-In this integral, we use a prior distribution for derived from our astrophysical prior.
-Note that the prior on can be very different for templates in different regions of parameter space.
-For example, the median of our prior for is 0.04 for Bank 0 (a low-mass bank), while it is 0.3 for Bank 14 (a high-mass bank).
-If we did not include such a prior, we would indeed expect to encounter a large loss of sensitivity on including HM.
-Comparing the performance of our search on events detected in previous -only catalogs (see Tables 3 and 6), we can confirm that our search sensitivity for these events has not been significantly degraded and is roughly comparable.
-We also find that in our new search, the significance of a few previously reported marginal events is increased significantly (GW190711_030756, GW200216_220804).
-Especially for the case of GW190711_030756, we suspect that the increase in sensitivity is due to addition of HM, but to test this we plan to do an injection study in the future.
-The downside of using a prior distribution on is that it is dependent on our choice of astrophysical prior (which is currently not precisely known).
-Especially, the mass ratio distribution is not very well constrained and affects our prior for .
-Using a different distribution for than the true astrophysical distribution can indeed lead to some loss of sensitivity.
-We chose a broad prior (flat in ) in our search, with the motivation that such a prior can be easily reweighted later in a hierarchical population analysis.
-Even though including HM can cause some loss in sensitivity for near-threshold equal mass systems, note that detecting even a few events with asymmetric-mass ratios can inform our population estimates much more than detecting a slightly larger number of equal-mass systems.
-It is worth drawing an analogy here to the case of spin distribution of templates used in -only searches.
-Neglecting templates with negative indeed improves the search sensitivity to near-threshold systems, but finding even a few systems can greatly help us distinguish between progenitor channels of BHs.
-We thus argue the goal of a search should not solely be to maximize the total number of detections, but to holistically cover most of the interesting regions of the binary parameter space.
-A similar argument also holds for a precessing search, where detecting even a few events with significantly misaligned spins can shed more light on the formation channels of binaries as compared to detecting a larger number of near-threshold aligned/zero spin mergers.
+Our sensitivity is higher than other searches especially for high mass and/or asymmetric mass ratio events due to addition of HM.
+Having higher sensitivity can improve over some regions of the population hyperparameter space even if we have null/marginal detections.
+We also find that in our new search, the significance of a few previously reported marginal events is increased significantly.
+Especially for the case of GW190711_030756, the increase in sensitivity is due to addition of HM.
+This will also be important in population level studies.
+Currently the LVK collaboration uses events with IFAR 1 year in their population studies.
+Due to the improved significance of our pipeline, more events can now be used in the IFAR 1 analysis
 VI Conclusions
-Nearly all gravitational searches performed till now only include the quadrupole mode of the full GW waveform predicted by general relativity.
+Nearly all the previous gravitational-wave searches only include the quadrupole mode of the full GW waveform predicted by general relativity.
 In our companion papers Wadekar et al.
-(2023a, b), we develop new efficient methods to introduce higher-order modes (HM) in our search pipeline (section II).
-This helped us perform the first search over the full binary black hole parameter space (i.e., over all binary inclinations) with including HM in our templates.
+(2023, 2024), we develop new efficient methods to introduce higher-order modes (HM) in our search pipeline (section II).
+This helped us perform the first search over the full binary black hole parameter space with including HM in our templates.
 Compared to our previous pipeline, we introduced additional improvements to remove/downweight noise transients (especially at low frequencies), which further improved our sensitivity to high mass and high redshift systems (section II.3).
-We searched over public data from the third LIGO–Virgo observing run and found 14 new candidate events with (see Table 1 and section III).
-Some of the new candidate events have several interesting astrophysical properties (section IV) such as: () masses in the IMBH range and pair instability mass gap, () high redshifts, () preference for positive effective spins, () asymmetric mass ratios.
-By simply summing over the complements of values of our new candidate events from Table 1, we expect that roughly 4 of the 14 new events are noise transients instead of being astrophysical signals.
-However, we note that estimates of and source parameters depend on the choice of prior, and we discuss this dependence in section V555We provide our posterior samples at https://github.com/JayWadekar/GW_higher_harmonics_search, which can be used to reweight our results for a different astrophysical prior..
-We also compared our pipeline’s results for the events published in previous IAS catalogs (Tables 2 & 6), and in the GWTC-3 and 4-OGC catalogs (Table 3).
-We broadly recovered the high-significance events, except some events which were either vetoed by our pipeline or fell below our trigger collection SNR threshold, which we discuss in detail in section III.3 (we also mention some possible future improvements in this regard).
-Our methodology for efficiently including realistic effects like HM in GW templates will continue to be useful for high-precision searches in upcoming and future gravitational wave experiments Fairhurst et al.
+We searched over public data from the third LIGO–Virgo observing run and found 11 new candidate events with (see Table 1 and section III).
+Some of the new candidate events have several interesting astrophysical properties (section IV) such as: () masses in the IMBH range and pair instability mass gap, () high redshifts, () non-zero effective spins, () asymmetric mass ratios.
+By simply summing over the complements of values of our new candidate events from Table 1, we expect that roughly 3 of the 11 new events are noise transients instead of being astrophysical signals.
+However, we note that estimates of and source parameters depend on the choice of prior, and we discuss this dependence in section V666See https://github.com/JayWadekar/GW_higher_harmonics_search for obtaining our posterior samples and the selection function of our pipeline..
+We also compared our pipeline’s results for the events published in previous IAS catalogs (Tables 2 & 6), and in the GWTC-3 and 4-OGC catalogs (Table LABEL:tab:lvc_events).
+We broadly recovered the high-significance events, except some events which were either vetoed by our pipeline, which we discuss in detail in section III.3.
+Our methodology for efficiently including realistic effects like HM in GW templates will continue to be useful for high-precision searches in upcoming and future gravitational wave experiments Ng et al.
+(2023); Fairhurst et al.
 (2023); Maggiore et al.
@@
 (2021).
-One particular example in which HM will be especially useful is when searching for extreme-mass-ratio inspirals in LISA Zhang et al.
+One particular example in which HM will be especially useful is when searching for intermediate to extreme mass-ratio inspirals Arun et al.
+(2007a, b); Porter and Cornish (2008); Berti et al.
 (2006); Baibhav et al.
+(2020); Pitte et al.
+(2023); Amaro-Seoane et al.
+(2023); Gair et al.
+(2008); we report results from such a search in a separate follow-up paper Ho-Yeuk Cheung et al.
 (2025).
-VI.1 Future work
-We plan to inject simulated gravitational-wave signals in the data and compare the fraction of injections recovered by our pipeline.
-The injections will help us determine the gain/loss in volume-time sensitivity in different regions of the parameter space due to including HM in our search.
-The injections will also help us gauge the level of improvement due to the additional glitch treatment methods that we described in section II.3.
-We plan to use the injection parameters from Abbott et al.
-(2021), which will help us perform an apples-to-apples comparison of our sensitivity with other pipelines.
-Having injections will also help quantify the selection function of our current pipeline and help in population inference studies.
-The new events reported here have moderate IFARs (the best event has yr), thus a methodology to combine confident and marginal events in population inference will be required Roulet et al.
-(2020).
-We expect our sensitivity to be higher than other searches for asymmetric mass ratio events due to addition of HM.
-Our sensitivity could also be higher for high-mass and high-redshift regions due to using improved glitch treatment methods and adding HM.
-Note that, in certain regions of the parameter space, having higher sensitivity can improve population estimates even if we have null/marginal detections.
-We note some interesting population inference questions in section IV.
-Currently, nearly all searches also ignore the effects of precession or eccentricity in their waveforms.
-Some of the challenges for including these effects are similar to the ones for HM: significant increase in the template bank size and loss of sensitivity due to collecting additional background.
-We hope to develop methods similar to the ones implemented in our search to help tackle those searches in the future.
+Note added– While this paper was on arXiv, a confident high-mass detection (GW231123) was reported by the LVK collaboration Abac et al.
+(2025) in the O4a data.
+Incorporating this detection into the search astrophysical prior could make our high-mass detections () more significant.
 Acknowledgements.
-We thank Horng Sheng Chia, Ajith Parameswaran, Will Farr, Vishal Baibhav, Katerina Chatziioannou, Vicky Kalogera, Sylvia Biscoveanu, Amanda Farah, Daniel Holz, Salvatore Vitale, Tom Edwards, Zoehyr Doctor, Suvodip Mukherjee and Mukesh Singh for helpful discussions.
+We thank Horng Sheng Chia, Ajith Parameswaran, Will Farr, Vishal Baibhav, Katerina Chatziioannou, Vicky Kalogera, Sylvia Biscoveanu, Amanda Farah, Daniel Holz, Salvatore Vitale, Tom Edwards, Zoheyr Doctor, Suvodip Mukherjee and Mukesh Singh for helpful discussions.
 DW gratefully acknowledges support from the Friends of the Institute for Advanced Study Membership and the Keck foundation.
+JR acknowledges support from the Sherman Fairchild Foundation.
 TV acknowledges support from NSF grants 2012086 and 2309360, the Alfred P.
@@
 Additional support for Advanced LIGO was provided by the Australian Research Council.
-Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain.Appendix A Additional plots and tables
+Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain.Appendix A Changes since previous arXiv version
+The previous version of our paper on the arXiv ([v1] 11 Dec 2023) included the list of events from an older version of our pipeline.
+Since that point, we have made a number of improvements to the pipeline.
+Furthermore, unlike the older version, the newer version of our IAS-HM pipeline has gone through an extensive injection recovery analysis to test its senstitivity (see section V.1).
+We list the details of other improvements below.
+We implemented a few extra noise reduction algorithms like TIER (see section II.3.1).
+We see an overall improvement in sensitivity due to it Wadekar et al.
+(a).
+We also however see both positive and negative fluctuations in IFAR and values of the near-threshold new candidate events.
+This is expected, as some of the near-threshold candidate events were probabilistically likely to be glitches and the TIER algorithm is able to downweight those.
+We fixed an issue with relative binning for template optimization in our bank BBH-1. This had earlier led to some of the LVK mergers being spuriously vetoed for this bank. We have now fixed the issue.
+In the previous version of our catalog, some of the fairly high SNR events in GWTC-3 with imprints of precession were vetoed by our pipeline.
+We added a ML-based veto overriding procedure (see section II.3.3) to override the veto for these events and we also find an improvement in the pipeline sensitivity due to this procedure.
+In section II.4, we discussed using an astrophysical prior for the templates.
+We improved the astrophysical prior applied to the templates using normalizing flows Wadekar et al.
+(b), superseding the earlier kernel density estimation based method given in Appendix C of O22.
+We also improved our coherent marginalized statistic using normalizing flows (further details will be provided in a separate paper Wadekar et al.
+(b)).
+Appendix B Additional plots and tables
 | Sr. | New candidate | This search | GWTC-3 sub-threshold | IAS 22-only sub-threshold | ||||||
 | No. | event | IFAR | IFAR | |||||||
-1 | GW190605_025957 | 75.7 | ||||||||
-2 | GW190806_033721 | 58.4 | ||||||||
-3 | GW190524_134109 | 60.0 | ||||||||
-4 | GW191113_103541 | 50.8 | Veto | |||||||
-5 | GW190615_030234 | 65.9 | ||||||||
-6 | GW190604_103812 | 46.8 | ||||||||
-7 | GW191228_195619 | 98.9 | Veto | |||||||
-8 | GW200304_172806 | 52.5 | ||||||||
-9 | GW190530_030659 | 60.0 | ||||||||
-10 | GW190511_163209 | 46.9 | ||||||||
-11 | GW200210_100022 | 53.7 | ||||||||
-12 | GW200301_211019 | 60.9 | ||||||||
-13 | GW190911_195101 | 65.7 | ||||||||
-14 | GW190530_133833 | 53.3 |
+| 1 | GW200104_184028 | 66.0 | ||||||||
+| 2 | GW190425_133124 | 67.6 | ||||||||
+| 3 | GW200319_022728 | 61.1 | ||||||||
+| 4 | GW190605_025957 | 63.2 | ||||||||
+| 5 | GW200114_020818 | 157 | ||||||||
+| 6 | GW190524_134109 | 56.8 | ||||||||
+| 7 | GW191213_204950 | 66.2 | ||||||||
+| 8 | GW190530_030659 | 63.2 | ||||||||
+| 9 | GW200210_035448 | 57.4 | ||||||||
+| 10 | GW190908_161547 | 58.2 | ||||||||
+| 11 | GW190710_150742 | 59.2 | ||||||||
 | (O3a) Name | Bank | Best fit template | IFAR (yr) | ||||||
-GW190711_030756 | 7,1 | ||||||||
-GW190707_083226 | 7,1 | ||||||||
-GW190818_232544 | 8,1 | ||||||||
-GW190906_054335 | 7,1 | ||||||||
-GW190814_192009 | 11,0 | ||||||||
-GW190821_124821 | 0,1 | ||||||||
-GW190704_104834 | 0,0 | ||||||||
-GW190920_113516 | 0,0 | ||||||||
-GW190910_012619 | 5,0 | ||||||||
+| GW190711_030756 | [6, 2] | ||||||||
+| GW190818_232544 | [8, 1] | ||||||||
+| GW190707_083226 | [7, 1] | ||||||||
+| GW190814_192009 | [10, 0] | ||||||||
+| GW190704_104834 | [0, 0] | ||||||||
+| GW190906_054335 | [6, 1] | ||||||||
+| GW190821_124821 | [0, 1] | ||||||||
+| GW190910_012619 | [4, 0] | ||||||||
 | (O3b) Name | Bank | IFAR (yr) | |||||||
-GW200109_195634 | 10,0 | ||||||||
-GW191228_085854 | 0,2 | ||||||||
-GW200316_235947 | 0,0 |
+| GW200109_195634 | [9, 0] | ||||||||
+| GW191117_023843 | [10, 0] | ||||||||
+| GW200210_100022 | [7, 1] | ||||||||
+| GW191228_085854 | [0, 2] | ||||||||
+| GW200316_235947 | [0, 0] | ||||||||
 | Candidate | Bank | Best fit template | IFAR (yr) | |||||||
 | per bank | overall | |||||||||
-GW190708_211916 | 8,1 | |||||||||
-GW190907_111633 | 12,0 | |||||||||
-GW191121_042622 | 13,0 | |||||||||
-GW190810_180039 | 6,2 | |||||||||
-GW200304_182240 | 11,0 | |||||||||
-GW200104_184028 | 8,1 | |||||||||
-GW191127_114537 | 13,0 | |||||||||
-GW190423_083253 | 10,0 | |||||||||
-GW200324_083227 | 7,1 | |||||||||
-GW200312_093903 | 11,0 | |||||||||
-GW200308_231646 | 8,1 | |||||||||
-GW190620_221103 | 7,1 | |||||||||
-GW200211_113417 | 11,0 | |||||||||
-GW200128_115507 | 8,1 | |||||||||
-GW190606_061230 | 0,2 | |||||||||
-GW190928_163629 | 4,2 | |||||||||
-GW200208_194804 | 13,0 | |||||||||
-GW200217_211137 | 9,0 | |||||||||
-GW190426_180646 | 7,1 | |||||||||
-GW200307_084747 | 9,0 | |||||||||
-GW200216_200717 | 8,1 | |||||||||
-GW190403_061805 | 6,0 | |||||||||
-GW200226_083409 | 6,2 | |||||||||
-GW190521_071322 | 7,1 | |||||||||
-GW190908_161547 | 7,1 | |||||||||
-GW190911_021724 | 4,2 | |||||||||
-GW200303_074125 | 7,1 | |||||||||
-GW190526_231941 | 7,1 | |||||||||
-GW200319_022728 | 8,1 | |||||||||
-GW200303_083614 | 0,1 | |||||||||
-GW191206_163102 | 13,0 | |||||||||
-GW191203_034221 | 3,1 | |||||||||
-GW200305_225311 | 3,1 | |||||||||
-GW200126_225208 | 7,1 | |||||||||
-GW191209_143052 | 0,1 | |||||||||
-GW200128_113935 | 4,2 | |||||||||
-GW200201_030733 | 4,2 | |||||||||
-GW200202_160037 | 4,2 | |||||||||
-GW191209_191159 | 14,0 | |||||||||
-GW200212_101822 | 0,3 | |||||||||
-GW190911_025827 | 14,0 | |||||||||
-GW200303_013439 | 0,2 | |||||||||
-GW190802_095334 | 5,2 | |||||||||
-GW191203_121111 | 14,0 | |||||||||
-GW190928_223455 | 2,0 | |||||||||
-GW191227_044656 | 14,0 |
+| GW200104_085935 | [10, 0] | |||||||||
+| GW190521_075726 | [14, 0] | |||||||||
+| GW190529_023648 | [9, 0] | |||||||||
+| GW190512_001054 | [6, 2] | |||||||||
+| GW191205_033000 | [7, 1] | |||||||||
+| GW200210_100022 | [7, 1] | |||||||||
+| GW200217_211137 | [9, 0] | |||||||||
+| GW191228_195439 | [10, 0] | |||||||||
+| GW191127_114537 | [12, 0] | |||||||||
+| GW190514_065259 | [3, 2] | |||||||||
+| GW191208_080334 | [5, 2] | |||||||||
+| GW191228_062343 | [11, 0] | |||||||||
+| GW190404_121150 | [9, 0] | |||||||||
+| GW200202_160037 | [5, 2] | |||||||||
+| GW190810_180039 | [5, 2] | |||||||||
+| GW190605_053717 | [1, 2] | |||||||||
+| GW200129_132219 | [5, 0] | |||||||||
+| GW200325_090608 | [10, 0] | |||||||||
+| GW200312_093903 | [11, 0] | |||||||||
+| GW200224_043639 | [12, 0] | |||||||||
+| GW200216_200717 | [7, 1] | |||||||||
+| GW190617_041509 | [8, 1] | |||||||||
+| GW191230_072040 | [5, 2] | |||||||||
+| GW200114_131740 | [0, 2] | |||||||||
+| GW191105_105458 | [6, 2] | |||||||||
+| GW191228_085854 | [0, 2] | |||||||||
+| GW190411_075756 | [9, 0] | |||||||||
+| GW191224_075833 | [10, 0] | |||||||||
+| GW200126_225208 | [6, 2] | |||||||||
+| GW200301_211019 | [2, 1] | |||||||||
+| GW200128_115507 | [8, 1] | |||||||||
+| GW190604_103812 | [11, 0] | |||||||||
+| GW190628_185055 | [1, 2] | |||||||||
+| GW190806_033721 | [9, 0] | |||||||||
+| GW190705_164632 | [7, 1] | |||||||||
+| GW191222_085357 | [8, 1] | |||||||||
+| GW190610_081435 | [1, 2] | |||||||||
+| GW191102_232120 | [2, 2] | |||||||||
+| GW200303_083614 | [0, 1] | |||||||||
+| GW200222_005735 | [0, 4] | |||||||||
+| GW191115_141720 | [9, 0] | |||||||||
+| GW200308_231646 | [8, 1] | |||||||||
+| GW190813_040521 | [5, 2] | |||||||||
+| GW190826_212543 | [8, 1] | |||||||||
+| GW190422_175505 | [8, 1] | |||||||||